### v2

# Reassessing Collaborative Writing Theories and Frameworks in the Age of LLMs: What Still Applies and What We Must Leave Behind

Abstract.
In this paper, we conduct a critical review of existing theories and frameworks on human-human collaborative writing to assess their relevance to the current human-AI paradigm in professional contexts, and draw seven insights along with design implications for human-AI collaborative writing tools. We found that, as LLMs nudge the writing process more towards an empirical “trial and error” process analogous to prototyping, the non-linear cognitive process of writing will stay the same, but more rigor will be required for revision methodologies. This shift would shed further light on the importance of coherence support, but the large language model (LLM)’s unprecedented semantic capabilities can bring novel approaches to this ongoing challenge. We argue that teamwork-related factors such as group awareness, consensus building and authorship – which have been central in human-human collaborative writing studies – should not apply to the human-AI paradigm due to excessive anthropomorphism. With the LLM’s text generation capabilities becoming essentially indistinguishable from human-written ones, we are entering an era where, for the first time in the history of computing, we are engaging in collaborative writing with AIs at workplaces on a daily basis. We aim to bring theoretical grounding and practical design guidance to the interaction designs of human-AI collaborative writing, with the goal of enhancing future human-AI writing software.
1. Introduction
As large language models (LLMs) increasingly become our daily co-writers in a uniquely asymmetrical relationship unlike any we have previously encountered, there is a need to reassess foundational theories and frameworks of collaborative writing to adapt to the human-AI paradigm. Their capabilities in natural language generation are becoming indistinguishable from human writing (Hämäläinen et al., 2023; Clark et al., 2021), but the mechanism of human-AI collaborative writing is clearly distinct from a human-human one. To reflect this paradigm shift, new frameworks and taxonomies are emerging, from domain-specific taxonomy development (Lee et al., 2024b), recognizing human-side effort through entropy and information gain (Wan et al., 2024), getting inspiration from the Overall Equipment Effectiveness Model (Wilbers et al., 2024), to exploring a non-linear framework to go beyond the linear human-only/human-AI/AI-only dimension of collaboration modes. (Hutson, 2025). While many of these proposals build on a partial aspect of existing theories and frameworks of collaborative writing, we believe that this must be approached holistically, synthesizing what past studies have taught us about the mechanisms of collaborative writing. It goes without saying that legacy theories and frameworks on collaborative writing have (naturally) focused only on the human-human mode. While the already established understandings of human-human collaborative writing should certainly apply to human-AI to some degree, we believe now is the moment to reassess our understanding of collaborative writing so far, and adapt it to the ever increasing human-AI mode.
Hence our research question: what insights can we draw from theories/frameworks on human-human collaborative writing to apply to the design of human-AI writing tools? Which parts of those theories/frameworks are not relevant for human-AI collaboration?
In this paper we conduct a critical review of legacy theories and frameworks on collaborative writing, and analyze what we see to be relevant to the human-AI paradigm as well as what only applies to the human-human paradigm, bringing clarity to the theoretical groundings required for the design of future human-AI collaborative writing tools. Our main findings are that, given the shift in the human’s role from hands-on writer to an editorial decision maker, the non-linear cognitive process of writing will stay the same, but more attention will be required for rigorous revision methodologies. In turn, this may reduce the focus on the planning aspects of collaborative writing, nudging writing more towards an empirical “trial and error” process than a fully planned out one, analogous to prototyping. This shift would require more focus on supporting coherence between human and AI writing, but we believe that the LLM’s unprecedented semantic capabilities can bring a break through to this ongoing challenge. Teamwork-related factors such as group awareness, consensus building and authorship, which have been central in human-human collaborative writing studies, will not apply to the human-AI paradigm due to excessive anthropomorphism, but it is equally the AI co-writer’s role to facilitate those factors between human writers to offload any unnecessary cognitive load. Finally, be it human-written or AI-written, writing will remain to be socially constructed by readers, which suggests that the consideration of reader profiles is inseparable from the text generation process. We provide practical design implications and example interface sketches with our insights to provide concrete design directions.
1.1. Study scope
The scope of this paper is limited to the pragmatic support of writing tasks in organizational workplace settings. This includes cases such as clinicians drafting discharge summaries, course coordinators drafting a response to student questions on online forums, meeting participants drafting a summary to send out to their team, or academic staff crafting grant proposals. We do not consider in fictional writing or rhetoric, nor do we consider topics around authorship, emotional satisfaction, or the deskilling aspects of writing. In particular, collaborative writing has traditionally had much focus on ownership and authorship, but building on Schneiderman (Shneiderman, 2022) we take the stance that LLMs are and should merely be a tool to help us get the job done in the most effective manner, and avoid any unnecessary forms of anthropomorphism in framing the AI as a “team mate”.
2. Background
We now revisit previous research situated in the social computing paradigm, collaborative writing, and human-AI collaboration, all of which inform our review.
2.1. Social Computing Paradigm
The Computers As Social Actors (CASA) paradigm (Nass et al., 1994), which revealed that humans apply the same social rules they would apply to humans to computers, has been a dominant framework in Human Computer Interaction (HCI). The underlining premise is simple: the way humans interact with computers is inherently social. The social response theory (Nass and Moon, 2000) took this a step further, explaining that such social rules are “mindlessly applied”, and that alternative explanations such as anthropomorphism simply cannot explain this phenomenon. Hence, to design interactions for new technologies, it is best to reframe computer interaction as a social interaction, and appropriately lean on relevant studies in social sciences along the way.
This approach of reframing a computer interaction as social interaction has been evident in various areas of AI studies. For example, in the field of Explainable AI, Miller (Miller, 2019) suggests that we build on what the social sciences have already uncovered about human-to-human explanation mechanisms, because after all, “explanations are social”, and being in a vacuum of computer science literature would not solve the problem.
Recently, such an approach has begun to become evident in dialogue studies, as represented in the emerging field of Human-Machine Communication (HMC), an area of communication research that regards technology as “communicative subjects, instead of mere interactive objects” (Guzman and Lewis, 2020). That is, HMC treats machines no longer as mediators, but as actors directly exchanging messages with humans. This reframing expands the focus of communication studies from human-human only to include human-AI contexts. For example, while it is well known that the lexical, syntactic and semantic alignment between two people engaged in a dialogue is crucial for mutual understanding (Pickering and Garrod, 2004), Ostrand and Berger claim that such alignment should be properly considered when building LLMs as well (Ostrand and Berger, 2024) for improved usability and personalization. This has also led to interesting discussions on the effects such alignment would have on bias amplification and echo chambers (Knoeferle et al., 2025).
With collaborative writing, however, there is still little work at the intersection of traditional human-human collaborative writing and the emerging human-AI writing paradigm. This is perhaps due to the fact that, until recently, writing with AI primarily referred to assistance tools such as spell- or grammar-checking. It also remains unclear as to what extent writing with LLMs can be considered “collaborative writing”. However, when a human tasks an LLM to write up a draft of a document, the human-AI pair is working “toward coordinated consensus that is reflected in a document that is written by one of the team members”, which we see to be undeniably qualifying as a form of collaborative writing (Lowry et al., 2004) (it is also worth noting that, according to Posner and Baecker (Posner and Baecker, 1992), single writer strategy was actually the most common strategy adopted even in human-only groups). If that is the case, the underlining reframing that is needed here remains the same. Writing is hardly a solitary act, be it with humans or with AI: it happens through the participation and interaction with others (Ede and Lunsford, 2006). We can see here that writing, too, is inherently social. As Wang et al. (Wang et al., 2020) suggest, it is crucial to integrate existing theories from HCI and Computer-Supported Cooperative Work (CSCW) in designing for the optimal form of human-AI collaborative writing.
2.2. Collaborative Writing
In 1969, Michel Foucault questioned the notion of the solitary author in the book “What is an author?”, stating that the author’s role is to regulate meaning and the circulation of text in an ever changing society (Foucault, 1979). The cognitive process theory of writing (Flower and Hayes, 1981), a widely adopted theory for unveiling an individual’s writing process, was updated by Hayes 15 years later (Hayes, 1996) to add more emphasis to the “social” aspect of writing, incorporating readers and collaborators into their framework. Ede and Lunsford (Lunsford and Ede, 1986) were the among the first to initiate collaborative writing as an academic research topic in 1986, reporting that 87 percent of respondents from six major professions sometimes wrote as a team or a group. Through further analysis they found that there were two distinct modes in collaborative writing: the “hierarchical” mode where one person has control and requires structure and division of labor, and the “dialogic” mode where all participants have equal input with shared authorship, which requires consensus and compromise (Ede and Lunsford, 2006). However, they also quickly realized that there was a significant lack of shared consensus on what collaborative writing is. As Allen et al. (Allen et al., 1987) explain: “very little detail is known about collaborative writing processes in general”. Due to the highly complex nature of document structures as well as the strong connection between writing and subjective emotions, collaborative writing is perhaps one of the most difficult forms of collaboration. As Farkas explains, there may not even be a common language for collaborative writing beyond a certain level (Lay and Karis, 1991). In light of this, Lowry et al. (Lowry et al., 2004) established the taxonomy of collaborative writing, outlining not just the various roles but also the writing strategies that groups adopt when writing together. As the role of computers and software becomes increasingly important in writing, this taxonomy is still getting updated to this day, including Collage (Buschek, 2024), CoCo matrix (Wan et al., 2024), Overall Writing Effectiveness (Wilbers et al., 2024), neurodiversity-inspired non-linear framework (Hutson, 2025), taxonomy of “text functions” (Larsen-Ledet et al., 2020) to name a few. Our work builds on these studies of emerging taxonomies, clarifying the limitations of existing human-human taxonomies and suggesting what needs to be renewed.
LLMs have brought a turning point to this field with their impressive text generation and semantic analysis capabilities. Given that writing tasks in educational and communication contexts are the second most frequent usage of Claude (Handa et al., 2025), an unprecedented number of workers are engaging in some form of collaborative writing with AIs on a daily basis. We believe it is more important now than ever to unveil the mechanisms of collaborative writing, especially between humans and AI.
2.3. Human-AI collaboration
Achieving synergy between human and AI performance has been the core focus of human-centered AI and human-robot interaction studies. One example is Explainable AI (XAI), attempting to achieve complementary performance between humans and AI for various tasks such as decision making and classification, through the transparency of the otherwise black box AI. Allowing users to drill down and ask follow up questions until they are satisfied has been an often discussed yet technically difficult challenge in XAI (Weld and Bansal, 2019), but LLMs have dramatically enabled us to study this approach further.
In general, however, the various approaches in XAI have led to mixed results with regards to its effectiveness on performance (Miller, 2023), which leads to Vaccaro et al.’s (Vaccaro et al., 2024) impressive meta-analysis suggesting that creation-related tasks (including writing) with GenAI statistically have more potential for achieving synergy– the positive complementary outcome from human-AI collaboration – and that the appropriate allocation of tasks between humans and AI may be the key to further progression. This is in line with past work (Zhang et al., 2020) which found that showing AI confidence or local feature explanations had no or detrimental effect to decision accuracy, suggesting that XAI can be most effective when the error boundaries of human’s and AI’s are as distinct as possible.
In this sense, we believe that the field of collaborative writing has much room for development when viewed from the human-AI collaboration lens. AI-assisted features such as auto-complete and the real-time collaboration capabilities of tools such as Google Docs or Overleaf have been studied broadly, but when it comes to collaborating with AI as a co-writer, we believe the appropriate understanding of their uniqueness and how they tie in to the already established understandings of human-human collaborative writing is yet to be defined. Similar to Yang et al. (Yang et al., 2022), we intentionally separate AI-assisted writing from AI as the co-writer.
Existing studies for AI as a co-writer have mostly focused on creative writing (Clark et al., 2018; Yang et al., 2019; Calderwood et al., 2020; Yang et al., 2022). Clark et al. (Clark et al., 2018) suggests that by reducing the intrusiveness of AI assistance for creative writing and adopting a pull method for the interactions, it can “allow people to write more closely to their normal writing process.” Yang et al. (Yang et al., 2022) took a turn-taking approach in building a “finish the story” system, finding that when users are equipped with the ability to edit AI generated text, they require only partial readability for the AI outputs because they can go in and fix the fluency of the overall text themselves. At the same time, in the adjacent field of image generation, Wadinambiarachchi et al. (Wadinambiarachchi et al., 2024) studied the effects that Midjourney111See https://www.midjourney.com/ – a generative AI system for image creation – has on non-designers and found that, without appropriate literacy on how to use the tool, GenAI support could lead to higher fixation of ideas, fewer ideas, less variety, and lower originality. This is a common challenge that is similarly seen in the field of Explainable AI, known as over-reliance (Vasconcelos et al., 2023). Due to the context-dependent nature of LLMs and GenAI, it is difficult to holistically understand the capabilities of LLM to apply to interaction designs. To tackle this, Lee et al. (Lee et al., 2022) proposes CoAuthor, a large dataset of recorded interactions between writers and GPT-3 across 1445 writing sessions, which has paved the way for HCI researchers to analyze human-AI writing cases in detail.
Our review builds on this fundamental Human-AI collaboration research to draw out the best of human and AI performance, specifically for writing tasks in organizational workplace settings, and in a way that is rooted in existing theories and frameworks. To the best of our knowledge, this is the first work to holistically reassess legacy collaborative writing theories and frameworks for the LLM era.
3. Survey Method
Our review followed the PRISMA methodology (Moher et al., 2009). The PRISMA diagram for our review is presented in Figure 1. Our research questions for the systematic review were as follows:
RQ1: What theories/frameworks exist for collaborative writing that can be applied to LLM interactions for practical writing tasks today? Which theories/frameworks are not relevant?
RQ2: What studies in HCI already build on these theories/frameworks when designing AI-assisted interactions?
RQ3: What design implications can we draw from these theories/frameworks to apply to current human-LLM interaction?
Drawing on the principal search system selections by Moher et al. (Moher et al., 2009), we selected 2 databases (Web of Science and the ACM Digital Library) and used the queries in Table 1. Because the selection of the ACM Digital Library was primarily to address RQ2, we added an additional filtering for LLM related papers, which was not necessary for Web of Science due to its broader, non-HCI-specific scope. Given that collaborative writing has been studied for decades, we did not limit by publication year.
| Database | Query |
| Web of Science | ”collaborative writing” AND (”theory” OR ”theories” OR ”framework” OR ”taxonomy” OR ”taxonomies”) |
| ACM Digital Library | ”collaborative writing” AND (”theory” OR ”theories” OR ”framework” OR ”taxonomy” OR ”taxonomies”) AND (”large language models” OR ”llm” OR ”generative ai” OR ”gen ai”) |
These queries resulted in a total of 3,067 entries (2,965 from Web of Science, 102 from ACM). After removing 7 duplicates, we applied the inclusion/exclusion criteria in Appendix A to the 3,060 entries based on the titles and abstracts. This left us with 65 entries – we lightly reviewed the full text, and screened with the predefined QA checklist in Appendix A. We applied a threshold of for the QA checklist (i.e. we only retained papers that achieved a full score of ), leaving us with 22 entries. Finally, we added 6 papers from backward referencing, resulting in a total of 28 papers for the full review. For each paper, we extracted the collaborative writing theory or framework that the work introduced or built on top of. The results of the data extraction are presented in Table LABEL:tab:papers.
| Study | Year | Theory/framework |
| Buschek (Buschek, 2024) | 2024 | Builds on the concept of literary collage |
| Chakrabarty et al. (Chakrabarty et al., 2024) | 2024 | Builds on the Cognitive Process Theory of Writing |
| Cress and Kimmerle (Cress and Kimmerle, 2023) | 2023 | Builds on theory of mind |
| De Silva (De Silva, 2007) | 2007 | Builds on narrative-based writing + Rhetorical Structure Theory |
| Dhillon et al. (Dhillon et al., 2024) | 2024 | Builds on scaffolding |
| Duin (Duin, 1990) | 1990 | Proposes the terms for collaborative writing, particularly for ways of working as a team |
| Ede and Lunsford (Ede and Lunsford, 2006) | 2006 | Proposes the hierarchical mode and dialogic mode of collaborative writing |
| Ens et al. (Ens et al., [n. d.]) | 2011 | Builds on the social constructivist theory and Ede and Lunsford’s dialogic/hierarchical mode of writing |
| Fiore and Wiltshire (Fiore and Wiltshire, 2016) | 2016 | Builds on distributed situation awareness (DSA) theory, interactive team cognition (ITC) theory, macrocognition in teams (MiTs) theory |
| Flower and Hayes (Flower and Hayes, 1981) | 1981 | Proposes the Cognitive Process Theory of Writing |
| Gero et al. (Gero et al., 2023) | 2023 | Proposes a taxonomy of support dynamics that writers seek |
| Hayes (Hayes, 1996) | 1996 | Updates the Cognitive Process Theory of Writing to include the social and physical aspects of writing |
| Laban et al. (Laban et al., 2024) | 2024 | Proposes a three-stage information verification framework (Warn-Verify-Audit) |
| Larsen-Ledet et al. (Larsen-Ledet et al., 2020) | 2020 | Buildds on Lowry et al.’s taxonomies of collaborative writing |
| Lay and Karis (Lay and Karis, 1991) | 1994 | Builds on the Bakhtin Circle |
| Lee et al. (Lee et al., 2024a) | 2024 | Proposes a design space for collaborative writing |
| Li et al. (Li et al., 2024) | 2024 | Applies user centered design framework for collaborative writing |
| Lowry et al. (Lowry et al., 2004) | 2004 | Proposes a taxonomy for collaborative writing |
| Lowry and Nunamaker (Lowry and Nunamaker, 2003) | 2003 | Proposes taxonomies for collaborative writing strategies |
| Mendoza-Chapa et al. (Mendoza-Chapa et al., 2002) | 2000 | Builds on Vertegaal’s framework on group awareness |
| Posner and Baecker (Posner and Baecker, 1992) | 1992 | Proposes a taxonomy of collaborative writing process |
| Prasetyo and Bandung (Prasetyo and Bandung, 2015) | 2015 | Builds on Lowry et al.’s taxonomies of collaborative writing |
| Rahhal et al. (Rahhal et al., 2007) | 2008 | Builds on Rhetorical Structure Theory |
| Ruan et al. (Ruan et al., 2024) | 2024 | Proposes a framework for joint analysis of collaborative document revision |
| Wan et al. (Wan et al., 2024) | 2024 | Builds on the Cognitive Process Theory of Writing and proposes CoCo matrix |
| Wiethof et al. (Wiethof et al., 2021) | 2021 | Builds on social response theory and Uncanny valley |
| Wilbers et al. (Wilbers et al., 2024) | 2024 | Proposes Overall Writing Effectiveness framework to assess human-LLM writing |
| Zhang et al. (Zhang et al., 2023) | 2023 | Builds on Toulmin’s argument model and the Cognitive Process Theory of Writing |
To ensure an objective process, two reviewers (the first and second author) took the same 15 randomly selected entries and applied the inclusion/exclusion criteria screening to check agreement levels. This sample screening not only revealed substantial agreement between the 2 reviewers (Cohen’s ), it also helped to add specificity to the exclusion criteria, which the first author applied to the rest of the entries. The first author conducted the full-text screening and the data extraction alone, with the second and third authors helping to resolve any conflicts and achieve consensus.
4. Results
4.1. Differences between humans and LLMs as a co-writer
Before we highlight our insights on legacy theories and frameworks, we first explain what we found to be unique about LLMs as a co-writer in the context of collaborative writing. They will be the important foundations in making our argument for what theories/frameworks apply to human-AI.
4.1.1. Always available for action and do not judge
As obvious as this may sound, LLMs need no rest. They are “always” available (except due to technical downtime, etc.), and there is no need for social hesitation or awkwardness when asking for help. As Gero et al. (Gero et al., 2023)’s work revealed, writers feel there is no need to work hard to establish or maintain a good relationship with AI agents, which frees them from the worry of being judged. While this may seem contrary to the CASA paradigm, Ede and Lunsford (Ede and Lunsford, 2006) interestingly point out that in the most ideal form of collaborative writing, you should “let everything show - the stupidities, the backsliding and illogicalities, and everything else that goes into writing a text - as well as to be open to those things in the other person, and to the other person’s strengths”. This suggests that, if used appropriately, the lack of social norm could paradoxically be what allows for exceptionally effective collaboration, be it for human-human or human-AI.
In other words, writers are able to ask from LLMs much more than what they would request from humans, without worrying about social norms. For example, Laban et al. (Laban et al., 2024) experimented with executable edits which allowed authors to edit faster than simple chat interactions. There is much to borrow from AI-powered code editors such as Cursor222See https://www.cursor.com or Github Copilot333See https://github.com/features/copilot in this aspect, where it is common for suggestions to be applied to the codebase with a click, beyond suggestions via chat.
4.1.2. No need for consensus building
Within the various definitions for collaborative writing we found in our review, some necessitate that all decisions be “made by consensus” (Lay and Karis, 1991). Ede and Lunsford (Ede and Lunsford, 2006) also propose the dialogic mode of writing (as opposed to the hierarchical mode), where all participants have equal input, shared authorship, and is suited for creativity, requiring consensus and compromise. Lowry (Lowry et al., 2004) labels this as “reactive writing”. Since AI tools have no opinions, values, or goals, we believe there is no need to consider building consensus with AI. However, there is value in human writers making their values and goals clear to an AI writing tool, and for human co-authors to do so with potential support from an AI writing tool, as we will discuss in our findings.
4.1.3. Distinguish content from instructions
We found multiple studies that distinguish diegetic (text as content) and non-diegetic (text as instructions/prompts) modes of interaction for LLMs (Buschek, 2024; Wan et al., 2024; Dang et al., 2023). Cress and Kimmerle’s work (Cress and Kimmerle, 2023) also mention the importance of distinguishing prompts for rhetorical space and prompts for content space. These studies suggest that, when writing with LLMs, the user always needs to keep in mind and be clear about whether they are referring to the writing itself or how to write the writing, which is not as evidently necessary with a human co-writer.
4.1.4. Distrust is encouraged
As meaningless as it may be to state, machines don’t care. Therefore, a certain level of distrust and skepticism towards AI as a co-writer is, ironically, encouraged. This is clearly different from human-human collaborative writing, where the establishment of trust and camaraderie is seen as a major component for success (Ens et al., [n. d.]; Lowry et al., 2004). We build on the approach of Cai et al. (Cai et al., 2019a) and Gu et al. (Gu et al., 2021) in creating clinical decision support tools: instead of waiting for AI to be good enough to be “fully automatable one day” (Gu et al., 2021), we should accept its “imperfection” and move towards creating tools with careful HCI considerations, which, in turn, would lead to better calibrated trust (Jacovi et al., 2021) and utility while experiencing less workload (Cai et al., 2019a).
4.2. Seven insights and design implications from legacy theories/frameworks
We now introduce the seven insights we drew from our review, along with design implications for future human-AI collaborative writing tools. The results are presented in Table 3.
| Theme | Design implication |
| Writing process and strategies are iterative, non-linear and interchangeable any time | Allow for edits or additional prompts at any given moment during the process of writing |
| Supporting planning has been the primary focus in collaborative writing tools | Equip users with granular revision control |
| Design novel methods to effectively prototype complex documents | |
| Group awareness is a key to success | Provide a moment for the writer to reflect on their expectations and needs from the AI co-writer |
| Coherence support is an ongoing challenge | Provide visual representations of the document’s semantic structure for continuous human-AI alignment |
| Writing is socially constructed by readers | Generate and review documents with reader profile awareness |
| Lack of clarity with authorship | Design data structures and visual cues to distinguish human-written text from AI-written ones |
| Provide built-in verification methods for AI-generated text | |
| Collaborative writing tools serve as external cognition | Make sure that the AI co-writer is contributing to both the production of text as well as the facilitation of the process itself |
4.2.1. Writing process and strategies are iterative, non-linear and interchangeable any time
Cognitively, writing processes are highly flexible and non-linear. This is one of the key insights of the Cognitive Process Theory of Writing (Flower and Hayes, 1981), which was the first to model the cognitive process of writing and has been referred to by many collaborative writing studies since. They proposed a hierarchically-organized model that primarily consists of 3 steps: (1) planning; (2) translating; and (3) reviewing, but writers “do not march through these processes in a simple 1, 2, 3 order” (Flower and Hayes, 1981). The writing process is essentially a constant iteration of creating and revising goals and sub-goals, and therefore key subprocesses such as idea generation, revision and evaluation can and should interrupt other processes at any given moment during the process of writing. More simply put: revision is not something that happens after writers do their writing. Rather, it happens “while in the midst of generating a text” (Faigley and Witte, 1981).
This non-linear and interchangeable nature of the writing process is a common theme across various theories and frameworks. Ede and Lunsford (Ede and Lunsford, 2006), when explaining their proposed model of hierarchical and dialogic modes of writing, equally stress the importance of resisting the urge to treat writing as a simple dichotomy between the two: “it is inherently mixed and paradoxical”. Even prior to Google Docs, collaborative writing tools have supported parallel-partitioned writing, allowing members to view and edit each other’s work at any given moment, which is known to increase teamwork effectiveness (Lowry and Nunamaker, 2003).
Design implication #1: Allow for edits or additional prompts at any given moment during the process of writing
Even if AI co-writers become the ones that generate most of the actual text, it is still vital for collaborative writing tools to respect and support the human’s cognitive processes. While we found studies that quite literally build on the 3 steps of the Cognitive Process Theory of Writing and incorporate its non-linearity (Chakrabarty et al., 2024; Zhang et al., 2023), we feel that a key design implication here is the importance of allowing the user to seamlessly revise and test new ideas at the same speed as when they are writing themselves. That is, the user should not be forced to wait for the LLM’s text generation to complete. Rather, the generation should always run in the background and be interruptable, so that the user can make edits or request for additional prompts at any given moment during the collaborative writing process. For instance, imagine an interface where the text generation is constantly run in the background, with no explicit “generate” button. We see this to be much more in line with the human’s non-linear and interchangeable cognitive process of writing, compared to a chat based interaction where the reviews, edits and imaginations are forced to take place in a linear manner.
4.2.2. Supporting planning has been the primary focus in collaborative writing tools
Traditionally, supporting planning has been the primary focus in CW tools (De Silva, 2007; Zhang et al., 2023; Lowry and Nunamaker, 2003). Prasetyo and Bandung discuss the planning features of their e-book writing tool, but features for reviewing hardly go beyond simple annotations and comments. This is understandable, since the translation phase is mostly the humans doing their work, and revision requires the humans to read each other’s work: they were both primarily human-centric tasks. However, particularly in cases where the LLMs take the role of writer, we question the necessity of such emphasis on planning. Indeed, in Chakrabarty et al.’s study on supporting professional writers with LLMs (Chakrabarty et al., 2024), the AI’s contributions with the lowest retention rate (i.e. not kept until the final document) were those related to high-level planning such as “Give me ideas for why a dog and its owner might become estranged”, while the highest retention (i.e. kept in the final document) were those related to translation such as “What’s a better name than Suzanne”. Perhaps this is reflecting one of the core differences between human and AI writers: AI does not have a stake in the written content, and therefore it is seemingly unfruitful to plan together. The strength of the AI co-writer lies not in the meticulous refinement of the message to convey, but simply in the faster and easier translation of ideas to words, as well as rigorous revision of formatting and errors. Interestingly, this shift is not entirely new in writing with technology: studies show that undergraduate writers planned less before writing when they used a word processor than with pen and paper (Hayes, 1996; Haas, 2013). If the shift from pen and paper to word processors reduced the (presumed) need for planning, then it is natural to assume that there will be bigger impacts with the shift from manual writing to LLMs. Now that the generation of text itself can be done effortlessly within seconds, we see planning to primarily emerge as a result of the human writer reflecting on the AI-generated drafts. It is not that the value of planning is diminishing, but that AI co-writers may encourage users to get something down on paper first, then revise and plan accordingly, nudging writing more towards an empirical “trial and error” process than a fully planned out one, similar to prototyping. Buschek’s work on comparing the current human-AI collaborative writing scene with literary collage from avant-garde literature (Buschek, 2024) frames this shift elegantly: writing with LLMs is closer to the assembling of “prefabricated text”, consisting of multiple voices and sources.
Design implication #2: Equip users with granular revision control & Design novel methods to effectively prototype complex documents
Shifting the focus from planning to revision, however, also comes with a cost when it comes to LLMs. LLMs are prone to sycophancy bias (Sharma et al., 2023), essentially making it a “yes man” regardless of factual truth. If so, is it also unrealistic to expect genuine critical feedback from LLMs? Hints can be drawn from the Re3 framework (Ruan et al., 2024), which proposes the taxonomy of revision in Figure 3. By dividing revision by its qualitative dimensions such as action, intent and granularity, and further subdividing each by their characteristics, we could enable users to provide more granular instructions to LLMs on the level of desired revision, such as requesting for grammar or clarity related edits for intent on sentence or subsentence granularity, but forbidding factual content updates which LLMs are notoriously known to be unreliable with. Given that the revision has long been mistakenly seen as just the “tidying-up activity” after completing a first draft (Faigley and Witte, 1981), helping the human writers to get accustomed to these more granular terminologies around revision is then perhaps the key agenda here.
We also acknowledge that the instantaneous and effortless text-generation can potentially lead to the loss of critical thinking for the human writers, as already seen with image generation tools (Wadinambiarachchi et al., 2024). If we are to blindly generate documents or emails to send to our peers without reviewing their output, then the results would certainly be catastrophic, both for our cognitive process, and for workplace collegiality and responsibility. The problem here is not the ease of text generation, but the difficulty in reviewing and iterating fully developed text. Perhaps the question we should be asking ourselves here is: how can we allow for effective prototyping of complex documents? For example, by leveraging the LLM’s semantic capabilities it is now possible for us to draft with smaller or summarized chunks of text, then move on to a fully fleshed document, similar to evolving prototypes from low-fidelity to high-fidelity. We see this to be a highly interesting and novel challenge that we welcome researchers to build further.
4.2.3. Group awareness is a key to success
Because human-human collaborative writing relies heavily on the ability to reach consensus, group awareness has been seen as a key to its success (Mendoza-Chapa et al., 2002). Group awareness in collaborative writing is not just about having awareness of what other members are doing during the collaboration process. Lowry et al. (Lowry et al., 2004) break down the types of group awareness for collaborative writing into informal awareness (knowing where the other members are working), group-structural awareness (knowing the roles and responsibilities of other members), social awareness (knowing the motivation and emotional state of other members) and workspace awareness (knowing what other members are up to in the software in use). Similarly, Duin (Duin, 1990) state the importance of sharing “terms” when writing collaboratively. This includes sharing each writer’s preferred writing process, commitment levels and skills prior to commencing the collaboration process, as well as deciding on when to stop revising and ending the collaboration session with appreciation of each other’s work.
Design implication #3: Provide a moment for writers to reflect on their expectations and needs from the AI co-writer
While a direct application of human-human group awareness to human-AI collaboration could be considered excessive anthropomorphism, we believe that defining the human’s expectation for AI tool’s roles and boundaries prior to the writing could help mitigate under- and over-reliance, as seen in the field of Explainable AI. For instance, Cai et al (Cai et al., 2019b) found that, in clinical contexts, clinicians wished to know the training data, the medical views and the design objectives of the AI, far before engaging in the explanations of their diagnosis recommendations. As long as it is in service of improving the user’s motivation towards the task at hand, it may be worthwhile to provide a quick onboarding to align on the user’s expectations on the AI tools’s roles and capabilities prior to commencing the writing. We see Wiethof et al. (Wiethof et al., 2021)’s work to be in line with this principle: they provided users with a written self-introduction by the AI agent prior to the writing, which led to “realistic expectations”. However, their implementation also came with the literal adoption of Nass and Moon’s Social Response Theory (Nass and Moon, 2000) such as providing social identities to the agent, which we consider unnecessary anthropomorphism. Similarly, Cress and Kimmerle (Cress and Kimmerle, 2023) suggest refining the human-human group awareness to suite the more asymmetrical human-AI relationship, and suggests the need for AIs to develop theory of mind, the cognitive ability to be aware of others’ mental states. A good balance here may be to simply provide a moment for the writer to individually reflect on what support they are seeking from the AI co-writer, both to nurture the self-awareness as well as to inform the AI tool of the user’s expectations and goals, but not introduce the AI co-writer as another team mate with its own values and writing styles.
When it comes to multiple human authors writing collaboratively with an AI co-writer, then it also becomes the AI co-writer’s role to facilitate team communication and at times resolve conflicts or consensus between the humans, which is a critical factor in the offloading of cognitive load in teamwork as we explain in 4.2.7.
4.2.4. Coherence support is an ongoing challenge
It is crucial to have syntactic and semantic consistency between the different parts of the documents. This has been an ongoing challenge in collaborative writing, particularly in the CSCW context (De Silva, 2007). Given that consensus and group awareness are key factors in human-human collaborative writing, it is understandable that coherence becomes a vulnerable factor that gets lost easily in the complexity of human collaboration, even more so with creative writing (Wan et al., 2024). Studies that tackled this in our review (De Silva, 2007; Rahhal et al., 2007) used the Rhetorical Structure Theory to semi-programatically point out semantic incoherence in documents. In Rhetorical Structure Theory, the document is first broken into non-overlapping segments, and the relationships between each segment are categorized into pre-defined labels. Some relationships resemble equal importance, while some resemble a hierarchical relation. Given that such semantic classification is one of the key strengths of LLMs, we believe that, if designed appropriately, AI co-writer tools could support collaborative writing without jeopardizing coherence much better than human team mates.
Design implication #4: Provide visual representations of the document’s semantic structure for continuous human-AI alignment
While appropriate prompting could potentially be a solution for achieving lexical and syntactic alignment, semantic alignment (i.e. consistency in the flow and document structure) would require more intricate design intervention, since it is less about the LLM’s writing capabilities and more about the human-AI alignment. Visual representation of the semantic document structure as seen in VISAR (Zhang et al., 2023) serves as a great inspiration here. Providing a visual cue of the current document structure would not just allow the user to be in constant alignment with the AI, but would essentially force the LLMs to engage in some form of “planning”, which would not otherwise be technically needed (Elkins, 2024) (unless specified to do so, such as with Chain-of-Thought prompting). Ultimately, it would enable the user to have an understanding of “how their text works” (Rahhal et al., 2007), which will be valuable regardless of the mode of collaboration.
4.2.5. Writing is socially constructed by readers
The notion that writing is socially constructed by the reader has been a key aspect of studies on writing and rhetoric in general. Bahktin has been a key figure here, stating that writing is a form of dialogue, and is a passive act that is socially interpreted by the readers in an “ever changing context” (Lay and Karis, 1991). Interestingly, when Hayes updated the original Cognitive Process Theory of Writing, they added the reader as a social environment of writing (Hayes, 1996) as well. While this notion has primarily been utilized to deepen our understanding of writing processes, we see some studies that have taken a rather literal approach to this. For example, Li et al. (Li et al., 2024) suggests incorporating readers into the writing process by applying the concept of “user-centered design” to writing towards a “reader-centered writing” process, envisioning systems where the text is constantly and automatically adapted to each reader profiles.
Design implication #5: Generate and review documents with reader profile awareness
If readers are a part of the writing, then it would mean that readers are inseparable from the LLM’s text generation process. A straightforward implementation would be to simply include the intended audience profile to the prompt, but we do not see this to be the solution. The fundamental implication here is that writers should always be conscious of the reader’s grammar, language and existing knowledge to maximize the ease of comprehension (Rahhal et al., 2007). One possible approach is to incorporate reader profile awareness into the revision process, where the AI co-writer is given the ability to assess the text as a pre-defined audience to check for any inconsistencies or lack of information from their perspective. The risk here would be excessive personalization, leading to bias amplification and social echo chambers, like we are starting to see in the adjacent field of human-LLM dialogue (Knoeferle et al., 2025).
4.2.6. Lack of clarity with authorship
The concept of authorship has been an ongoing debate in collaborative writing (Ede and Lunsford, 2001). While we excluded papers that primarily focused on authorship during our screening, it has still come up as an evident theme. For instance, Dhillon et al. (Dhillon et al., 2024) explored the effects of scaffolding during human-AI collaborative writing, and found that high scaffolding (i.e. next paragraph suggestion) resulted in higher writing quality and productivity, but also resulted in the decrease of text ownership and overall satisfaction. As more and more writing is conducted by AI, it is inevitable that our sense of ownership towards the text will decrease. Since our scope for this paper is the pragmatic support of task completion, we do not necessarily see this as a problem. Having said that, the concept of authorship is directly tied with the author’s accountability (Stokel-Walker, 2023), and given that only the human writers can be kept accountable for the consequences of any human-AI collaborative writing (Stokel-Walker, 2023), design considerations are necessary to clarify where the accountability lies.
Design implication #6: Design data structures and visual cues to distinguish human-written text from AI-written ones & Provide built-in verification methods for AI-generated text
We believe that keeping track of human contribution vs LLM contribution is a crucial aspect in future collaborative writing tools. Laban et al’s (Laban et al., 2024) InkSync proposes an insightful JSON data structure that can constantly keep track of text revisions as well as flagging for additional fact checking. Such data structures would allow for collaborative writing tools to visually distinguish human written text from AI written ones, enhancing the notion of accountability for the user. Their work also proposes the Warn-Verify-Audit framework, where new information introduced to the document by LLMs are flagged, and can be automatically searched online or sent to an external auditor for manual fact checking. We see similar interesting fact-checking approaches emerging (Liu et al., 2024; Amirizaniani et al., 2024; Dietz, 2024), all of which are human-in-the-loop approaches to maximize the strengths of manual oversight and LLM automation. While these studies are beyond the scope of this paper, they would inevitably need to be eventually built into human-AI collaborative writing tools.
4.2.7. Collaborative writing tools serve as external cognition
Fiore and Wiltshire (Fiore and Wiltshire, 2016) explored the various theories related to technology as an artifact of external cognition to make the claim that tools are inseparable from a team’s cognition. This allows us to circle back to the very reason why human-AI collaboration is beneficial in work-context writing tasks in the first place: the primary purpose of involving LLMs in collaborative writing should be to get the job done in the most effective manner, and the offloading of cognitive resources is a key criteria in achieving this goal. As we design AI co-writer tools, we must take great care in making sure that the tool is not resulting in unwanted cognitive load, bringing more disruption than improvement. Unnecessary cognitive load can be witnessed at various moments throughout the writing process. For example, Dhillon et al. (Dhillon et al., 2024) found that low-level scaffolding (i.e. next-sentence suggestions) did not enhance writing productivity, due to its intended assistance being too “rudimentary and disruptive” for the user’s writing process. Instead, high-level scaffolding (i.e. next-paragraph suggestions) resulted in significant improvements in the writing quality, due to the effective reduction of cognitive load. Larsen-Ledet et al. (Larsen-Ledet et al., 2020) shed light from a socio-technical perspective, finding that writers often switch between multiple writing tools, not just for functional reasons, but to get away from the collaborative mode to “sit for myself.” That is, writers sometimes prefer to stay away from collaboration to stay focused.
Traditionally, cognitive load has been measured in HCI through post-task questionnaires, such as with the NASA task load index (Hart and Staveland, 1988), although recently the validated of this for HCI has been questioned (Babaei et al., 2025). Emerging frameworks such as the Overall Writing Effectiveness framework (Wilbers et al., 2024) could also serve as an alternative metric in surfacing the impacts that LLMs bring to collaborative writing.
Design implication #7: Make sure that the AI co-writer is contributing to both the production of text as well as the facilitation of the process itself
When working collaboratively, it is not just the task that becomes the source of cognitive effort, but also the process itself, implying that the AI co-writer’s role is not limited to producing high quality text, but also to make sure that the process itself is not causing discomfort for the user. This is inspired greatly by the way Fiore and Wiltshire (Fiore and Wiltshire, 2016) emphasize the importance of distinguishing teamwork (behaviors required for working together with others) and taskwork (functions required for getting the task done). Tools support teamwork through the visualization of group awareness, team progress and facilitating communication, while tools also support taskwork through the direct assistance of the task (e.g. writing or decision making). In other words, collaborative writing tools can help to offload cognitive load not just through the direct contribution to the writing as a co-writer, but also by streamlining the collaboration process itself, both for human-AI teamwork as well as human-human teamwork.
5. Discussion
We now come back to our initial research question: what insights can we draw from theories/frameworks on human-human collaborative writing to apply to the design of human-AI writing tools? Which parts of those theories/frameworks are not relevant for human-AI collaboration? We outline our views in Figure 10.
While the role of humans undeniably changes from a hands-on writer to more of an editorial decision maker (Buschek, 2024), we believe that the cognitive process of writing will not and should not change — at least for good writing. This is to make sure that the writer continues to stay accountable for the writing they produce (which the AI cannot). Because of this, the non-linear, interchangeable nature of writing that studies based on the Cognitive Process Theory of Writing (Flower and Hayes, 1981) will stay as a core foundation of future collaborative writing tool designs. However, because the speed and ease of translating ideas to words will increase dramatically, we believe the writing process to become closer to an empirical “trial and error” process analogous to prototyping, resulting in less need for fully fledged planning, and more support needed for rigorous revision methodologies. In particular, tools will need to equip writers with enhanced understanding and terminologies for different revision types, to draw more granular boundaries for the AI support.
This shift in turn would further elevate the importance of coherence support in human-AI collaborative writing, which has been an ongoing challenge in this field to this day. Given the unprecedented semantic classification capabilities of LLMs, we are optimistic about the impact that LLMs can bring to the alignment of document structures through real-time visualization, which can positively impact not just the human-AI collaborative writing but human-human writing as well.
Since we take the stance that AIs are no more than a tool and humans should always remain in control, teamwork-related factors such as group awareness, consensus building and authorship, which have been much discussed in collaborative writing thus far, will not apply in the human-AI paradigm, except between multiple human co-authors. Therefore, when it comes to the writing strategies, it seems inevitable that most human-AI collaborative writing will end up as single-author writing, or will be conducted in the hierarchical mode where the AI co-writer will primarily translate ideas to words instantaneously under human supervision, and the humans will primarily take an editorial role. The more equally distributed forms of collaborative writing, known as reactive writing, parallel writing or the dialogic mode, will remain a strategy unique to human collaboration. However, given that the reduction of cognitive load results not just from the direct support of text production but in streamlining the collaboration process as well, it is an equally important role of the AI co-writer to remove any unnecessary frictions in the collaboration process itself, particularly when there are more than one human writers involved.
Finally, the notion that writing is socially constructed through readers remains to be just as valid, and we urge researchers to explore the implications that LLM’s diverse text generation capabilities will have to this space. At the same time, we acknowledge that this will only apply while humans are the readers: if human and AIs are co-writing, there is no reason to doubt that humans and AIs will co-read, and if (willingly or unwillingly) documents are to be agentically processed in the future, this will signal another significant turning point for writing in general.
6. Limitations
Finally, we would like to acknowledge some general limitations of our literature analysis approach.
First, we chose to look at Web of Science and the ACM Digital Library, but it is possible that we are missing collaborative writing theories or frameworks that were not covered by these two databases. We are, however, confident that we have sufficiently covered at least the most influential ones commonly found in past literature.
Second, our findings apply only to the scope of collaborative writing that the authors have chosen to study, as explained in Section 1.1. Our results will not be applicable to collaborative writing in other contexts, such as fictional writing or educational settings – these may be explored in future work.
Third, we intentionally cast a wide net for the systematic review, trying to capture as many related theories and frameworks from as many disciplines as possible. This has led to the decision of not screening based on their impact or citations, which may have led to variations in the quality of the surveyed papers.
7. Conclusion
In this paper, we have argued for the need to reassess foundational theories and frameworks of collaborative writing to adapt to the human-AI paradigm, as LLMs increasingly become our daily co-writers in a uniquely asymmetrical relationship unlike any we have previously encountered. We conducted a critical review of existing theories and frameworks on human-human collaborative writing to assess their relevance to the human-AI paradigm, and drew seven insights and implications that can be applied to the design of human-AI writing tools. We found that, as LLMs nudge the writing process more towards an empirical “trial and error” process analogous to prototyping, the non-linear cognitive process of writing will stay the same, but more attention will be required for rigorous revision methodologies. In turn, this may reduce the focus on the planning aspects of collaborative writing, further increasing the importance of coherence support between human and AI writing. We believe that the LLM’s unprecedented semantic capabilities can bring a break through to this ongoing challenge, through real time visualizations of the document’s semantic structure. Teamwork-related factors such as group awareness, consensus building and authorship, which have been central discussions in human-human collaborative writing, will not apply to the human-AI paradigm, resulting in the majority of human-AI collaborative writing being categorized in the single-author writing strategy. Finally, writing will continue to be socially constructed by readers, but this is dependent on the fact that humans will continue to be the primary readers of the documents we create – which may not necessarily be the case in the near future. Beyond theoretical insights, we provide practical design implications and example interface sketches to guide real-world application. We hope that our work will bring clarity and theoretical grounding to the interaction designs of human-AI collaborative writing.
Acknowledgements.
The interface designs in this paper used Google’s Material Icons444See https://fonts.google.com/icons provided under the Apache License, Version 2.0.References
- (1)
- Allen et al. (1987) Nancy Allen, Dianne Atkinson, Meg Morgan, Teresa Moore, and Craig Snow. 1987. What experienced collaborators say about collaborative writing. Iowa State J. Bus. Tech. Commun. 1, 2 (Sept. 1987), 70–90.
- Amirizaniani et al. (2024) Maryam Amirizaniani, Jihan Yao, Adrian Lavergne, Elizabeth Snell Okada, Aman Chadha, Tanya Roosta, and Chirag Shah. 2024. LLMAuditor: A framework for auditing Large Language Models using human-in-the-loop. arXiv [cs.AI] (Feb. 2024).
- Babaei et al. (2025) Ebrahim Babaei, Tilman Dingler, Benjamin Tag, and Eduardo Velloso. 2025. Should we use the NASA-TLX in HCI? A review of theoretical and methodological issues around Mental Workload Measurement. International journal of human-computer studies 201, 103515 (June 2025), 103515. doi:10.1016/j.ijhcs.2025.103515
- Buschek (2024) Daniel Buschek. 2024. Collage is the new writing: Exploring the fragmentation of text and user interfaces in AI tools. arXiv [cs.HC] (May 2024).
- Cai et al. (2019a) Carrie J Cai, Emily Reif, Narayan Hegde, Jason Hipp, Been Kim, Daniel Smilkov, Martin Wattenberg, Fernanda Viegas, Greg S Corrado, Martin C Stumpe, and Michael Terry. 2019a. Human-centered tools for coping with imperfect algorithms during medical decision-making. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems. ACM, New York, NY, USA.
- Cai et al. (2019b) Carrie J Cai, Samantha Winter, David Steiner, Lauren Wilcox, and Michael Terry. 2019b. “hello AI”: Uncovering the onboarding needs of medical practitioners for human-AI collaborative decision-making. Proc. ACM Hum. Comput. Interact. 3, CSCW (Nov. 2019), 1–24.
- Calderwood et al. (2020) Alex Calderwood, Vivian Qiu, K Gero, and Lydia B Chilton. 2020. How novelists use generative language models: An exploratory user study. HAI-GEN+user2agent@IUI (2020).
- Chakrabarty et al. (2024) Tuhin Chakrabarty, Vishakh Padmakumar, Faeze Brahman, and Smaranda Muresan. 2024. Creativity support in the age of large language models: An empirical study involving professional writers. In Creativity and Cognition. ACM, New York, NY, USA.
- Clark et al. (2021) Elizabeth Clark, Tal August, Sofia Serrano, Nikita Haduong, Suchin Gururangan, and Noah A Smith. 2021. All that’s ‘human’ is not gold: Evaluating human evaluation of generated text. In Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing (Volume 1: Long Papers). Association for Computational Linguistics, Stroudsburg, PA, USA, 7282–7296.
- Clark et al. (2018) Elizabeth Clark, Anne Spencer Ross, Chenhao Tan, Yangfeng Ji, and Noah A Smith. 2018. Creative writing with a machine in the loop: Case studies on slogans and stories. In 23rd International Conference on Intelligent User Interfaces. ACM, New York, NY, USA.
- Cress and Kimmerle (2023) Ulrike Cress and Joachim Kimmerle. 2023. Co-constructing knowledge with generative AI tools: Reflections from a CSCL perspective. Int. J. Comput. Support. Collab. Learn. 18, 4 (Oct. 2023), 607–614.
- Dang et al. (2023) Hai Dang, Sven Goller, Florian Lehmann, and Daniel Buschek. 2023. Choice over control: How users write with large language models using diegetic and non-diegetic prompting. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. ACM, New York, NY, USA, 1–17.
- De Silva (2007) N H De Silva. 2007. Using narratives in collaborative writing: An example.
- Dhillon et al. (2024) Paramveer S Dhillon, Somayeh Molaei, Jiaqi Li, Maximilian Golub, Shaochun Zheng, and Lionel Peter Robert. 2024. Shaping human-AI collaboration: Varied scaffolding levels in co-writing with language models. In Proceedings of the CHI Conference on Human Factors in Computing Systems, Vol. 22. ACM, New York, NY, USA, 1–18.
- Dietz (2024) Laura Dietz. 2024. A workbench for autograding retrieve/generate systems. In Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval, Vol. 67. ACM, New York, NY, USA, 1963–1972.
- Duin (1990) Ann Hill Duin. 1990. TERMS AND TOOLS: A theory and research-based approach to collaborative writing. Bull. Assoc. Bus. Commun. 53, 2 (June 1990), 45–50.
- Ede and Lunsford (2001) Lisa Ede and Andrea A Lunsford. 2001. Collaboration and concepts of authorship. Publ. Mod. Lang. Assoc. Am. 116, 2 (March 2001), 354–369.
- Ede and Lunsford (2006) L S Ede and A A Lunsford. 2006. Singular texts/plural authors: Perspectives on collaborative writing. Southern Illinois University Press, Carbondale, IL.
- Elkins (2024) Katherine Elkins. 2024. AI comes for the author. Poet. Today 45, 2 (June 2024), 267–274.
- Ens et al. ([n. d.]) Anita H Ens, Karen Boyd, L Allyson Matczuk, and Warren T Nickerson. [n. d.]. Graduate Students’ Evolving Perceptions of Writing Collaboratively. Canadian Journal of Higher Education 41, 2 ([n. d.]), 62–81.
- Faigley and Witte (1981) Lester Faigley and Stephen Witte. 1981. Analyzing Revision. Coll. Compos. Commun. 32, 4 (Dec. 1981), 400.
- Fiore and Wiltshire (2016) Stephen M Fiore and Travis J Wiltshire. 2016. Technology as teammate: Examining the role of external cognition in support of team cognitive processes. Front. Psychol. 7 (Oct. 2016), 1531.
- Flower and Hayes (1981) Linda Flower and John R Hayes. 1981. A cognitive process theory of writing. Coll. Compos. Commun. 32, 4 (Dec. 1981), 365.
- Foucault (1979) Michel Foucault. 1979. Authorship: What is an Author? Screen 20, 1 (03 1979), 13–34.
- Gero et al. (2023) Katy Ilonka Gero, Tao Long, and Lydia B Chilton. 2023. Social dynamics of AI support in creative writing. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. ACM, New York, NY, USA, 1–15.
- Gu et al. (2021) Hongyan Gu, Jingbin Huang, Lauren Hung, and Xiang ’anthony’ Chen. 2021. Lessons Learned from Designing an AI-Enabled Diagnosis Tool for Pathologists. Proc. ACM Hum.-Comput. Interact. 5, CSCW1 (April 2021), 1–25.
- Guzman and Lewis (2020) Andrea L Guzman and Seth C Lewis. 2020. Artificial intelligence and communication: A Human–Machine Communication research agenda. New Media Soc. 22, 1 (Jan. 2020), 70–86.
- Haas (2013) Christina Haas. 2013. Writing technology: Studies on the materiality of literacy. Routledge.
- Handa et al. (2025) Kunal Handa, Alex Tamkin, Miles McCain, Saffron Huang, Esin Durmus, Sarah Heck, Jared Mueller, Jerry Hong, Stuart Ritchie, Tim Belonax, Kevin K Troy, Dario Amodei, Jared Kaplan, Jack Clark, and Deep Ganguli. 2025. Which economic tasks are performed with AI? Evidence from millions of Claude conversations. arXiv [cs.CY] (Feb. 2025).
- Hart and Staveland (1988) Sandra G Hart and Lowell E Staveland. 1988. Development of NASA-TLX (task load index): Results of empirical and theoretical research. In Advances in Psychology. Advances in psychology, Vol. 52. Elsevier, 139–183.
- Hayes (1996) J Hayes. 1996. A new framework for understanding cognition and affect in writing. The science of writing: Theories, methods, individual differences, and applications. 432 (1996), 1–27.
- Hutson (2025) James Hutson. 2025. Human-AI collaboration in writing: A multidimensional framework for creative and intellectual authorship. International Journal of Changes in Education (Feb. 2025).
- Hämäläinen et al. (2023) Perttu Hämäläinen, Mikke Tavast, and Anton Kunnari. 2023. Evaluating large language models in generating synthetic HCI research data: A case study. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems, Vol. 27. ACM, New York, NY, USA, 1–19.
- Jacovi et al. (2021) A Jacovi, A Marasović, T Miller, and others. 2021. Formalizing trust in artificial intelligence: Prerequisites, causes and goals of human trust in AI. Proceedings of the 2021 (2021). https://dl.acm.org/doi/abs/10.1145/3442188.3445923
- Knoeferle et al. (2025) Pia Knoeferle, Sebastian Möller, Dorothea Kolossa, Veronika Solopova, and Georg Rehm. 2025. How desirable is alignment between LLMs and linguistically diverse human users? arXiv [cs.CL] (Feb. 2025).
- Laban et al. (2024) Philippe Laban, Jesse Vig, Marti Hearst, Caiming Xiong, and Chien-Sheng Wu. 2024. Beyond the chat: Executable and verifiable text-editing with LLMs. In Proceedings of the 37th Annual ACM Symposium on User Interface Software and Technology, Vol. 10. ACM, New York, NY, USA, 1–23.
- Larsen-Ledet et al. (2020) Ida Larsen-Ledet, Henrik Korsgaard, and Susanne Bødker. 2020. Collaborative writing across multiple artifact ecologies. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems. ACM, New York, NY, USA.
- Lay and Karis (1991) Mary M Lay and William M Karis. 1991. Collaborative Writing in Industry: Investigations in Theory and Practice. Baywood Press.
- Lee et al. (2024a) Mina Lee, Katy Ilonka Gero, John Joon Young Chung, Simon Buckingham Shum, Vipul Raheja, Hua Shen, Subhashini Venugopalan, Thiemo Wambsganss, David Zhou, Emad A Alghamdi, Tal August, Avinash Bhat, Madiha Zahrah Choksi, Senjuti Dutta, Jin L C Guo, Md Naimul Hoque, Yewon Kim, Simon Knight, Seyed Parsa Neshaei, Antonette Shibani, Disha Shrivastava, Lila Shroff, Agnia Sergeyuk, Jessi Stark, Sarah Sterman, Sitong Wang, Antoine Bosselut, Daniel Buschek, Joseph Chee Chang, Sherol Chen, Max Kreminski, Joonsuk Park, Roy Pea, Eugenia Ha Rim Rho, Zejiang Shen, and Pao Siangliulue. 2024a. A Design Space for Intelligent and Interactive Writing Assistants. In Proceedings of the CHI Conference on Human Factors in Computing Systems, Vol. 7. ACM, New York, NY, USA, 1–35.
- Lee et al. (2024b) Minhwa Lee, Zae Myung Kim, Vivek Khetan, and Dongyeop Kang. 2024b. Human-AI collaborative taxonomy construction: A case study in profession-specific writing assistants. In Proceedings of the Third Workshop on Intelligent and Interactive Writing Assistants. ACM, New York, NY, USA, 51–57.
- Lee et al. (2022) Mina Lee, Percy Liang, and Qian Yang. 2022. CoAuthor: Designing a human-AI collaborative writing dataset for exploring language model capabilities. In CHI Conference on Human Factors in Computing Systems. ACM, New York, NY, USA, 1–19.
- Li et al. (2024) Ge Li, Danai Vachtsevanou, Jérémy Lemée, Simon Mayer, and Jannis Strecker. 2024. Reader-aware writing assistance through reader profiles. In Proceedings of the 35th ACM Conference on Hypertext and Social Media, Vol. 59. ACM, New York, NY, USA, 344–350.
- Liu et al. (2024) Houjiang Liu, Anubrata Das, Alexander Boltz, Didi Zhou, Daisy Pinaroc, Matthew Lease, and Min Kyung Lee. 2024. Human-centered NLP Fact-checking: Co-Designing with Fact-checkers using Matchmaking for AI. Proc. ACM Hum. Comput. Interact. 8, CSCW2 (Nov. 2024), 1–44.
- Lowry et al. (2004) Paul B Lowry, Aaron Curtis, and Michelle René Lowry. 2004. Building a taxonomy and nomenclature of collaborative writing to improve interdisciplinary research and practice. J. Bus. Commun. 41, 1 (Jan. 2004), 66–99.
- Lowry and Nunamaker (2003) P B Lowry and J F Nunamaker. 2003. Using internet-based, distributed collaborative writing tools to improve coordination and group awareness in writing teams. IEEE Trans. Prof. Commun. 46, 4 (Dec. 2003), 277–297.
- Lunsford and Ede (1986) A Lunsford and Lisa S Ede. 1986. Why write…Together: A research update. Rhetoric Review 5 (Sept. 1986), 71–77.
- Mendoza-Chapa et al. (2002) S Mendoza-Chapa, M Romero-Salcedo, and H Oktaba. 2002. Group awareness support in collaborative writing systems. In Proceedings Sixth International Workshop on Groupware. CRIWG 2000. IEEE Comput. Soc, 112–118.
- Miller (2019) Tim Miller. 2019. Explanation in artificial intelligence: Insights from the social sciences. Artif. Intell. 267 (Feb. 2019), 1–38.
- Miller (2023) Tim Miller. 2023. Explainable AI is Dead, Long Live Explainable AI! Hypothesis-driven Decision Support using Evaluative AI. In Proceedings of the 2023 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’23). Association for Computing Machinery, New York, NY, USA, 333–342. doi:10.1145/3593013.3594001
- Moher et al. (2009) David Moher, Alessandro Liberati, Jennifer Tetzlaff, Douglas G Altman, and PRISMA Group. 2009. Preferred reporting items for systematic reviews and meta-analyses: the PRISMA statement. BMJ 339, jul21 1 (July 2009), b2535.
- Nass and Moon (2000) Clifford Nass and Youngme Moon. 2000. Machines and mindlessness: Social responses to computers. J. Soc. Issues 56, 1 (Jan. 2000), 81–103.
- Nass et al. (1994) Clifford Nass, Jonathan Steuer, and Ellen R Tauber. 1994. Computers are social actors. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems. ACM, New York, NY, USA, 72–78.
- Ostrand and Berger (2024) Rachel Ostrand and Sara E Berger. 2024. Humans linguistically align to their conversational partners, and language models should too. (July 2024).
- Pickering and Garrod (2004) Martin J Pickering and Simon Garrod. 2004. The interactive-alignment model: Developments and refinements. Behav. Brain Sci. 27, 2 (April 2004), 212–225.
- Posner and Baecker (1992) I R Posner and R M Baecker. 1992. How people write together (groupware). In Proceedings of the Twenty-Fifth Hawaii International Conference on System Sciences, Vol. iv. IEEE, 127–138 vol.4.
- Prasetyo and Bandung (2015) Muhammad Eko Budi Prasetyo and Yoanes Bandung. 2015. Development of real-time collaboration system for e-book writing. In 2015 International Conference on Information Technology Systems and Innovation (ICITSI). IEEE, 1–7.
- Rahhal et al. (2007) Charbel Rahhal, Hala Skaf-Molli, Pascal Molli, and Nishadi De Silva. 2007. SemCW: Semantic collaborative writing using RST. In 2007 International Conference on Collaborative Computing: Networking, Applications and Worksharing (CollaborateCom 2007). IEEE, 484–493.
- Ruan et al. (2024) Qian Ruan, Ilia Kuznetsov, and Iryna Gurevych. 2024. Re3: A holistic framework and dataset for modeling collaborative document revision. In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers). Association for Computational Linguistics, Stroudsburg, PA, USA, 4635–4655.
- Sharma et al. (2023) Mrinank Sharma, Meg Tong, Tomasz Korbak, David Duvenaud, Amanda Askell, Samuel R Bowman, Esin Durmus, Zac Hatfield-Dodds, Scott R Johnston, Shauna M Kravec, Timothy Maxwell, Sam McCandlish, Kamal Ndousse, Oliver Rausch, Nicholas Schiefer, Da Yan, Miranda Zhang, and Ethan Perez. 2023. Towards Understanding Sycophancy in Language Models. In The Twelfth International Conference on Learning Representations.
- Shneiderman (2022) Ben Shneiderman. 2022. Human-Centered AI. Oxford University Press UK.
- Stokel-Walker (2023) Chris Stokel-Walker. 2023. ChatGPT listed as author on research papers: many scientists disapprove. Nature 613, 7945 (Jan. 2023), 620–621.
- Vaccaro et al. (2024) Michelle Vaccaro, Abdullah Almaatouq, and Thomas Malone. 2024. When combinations of humans and AI are useful: A systematic review and meta-analysis. Nat. Hum. Behav. (Oct. 2024), 1–11.
- Vasconcelos et al. (2023) Helena Vasconcelos, Matthew Jörke, Madeleine Grunde-McLaughlin, Tobias Gerstenberg, Michael S Bernstein, and Ranjay Krishna. 2023. Explanations can reduce overreliance on AI systems during decision-making. Proc. ACM Hum. Comput. Interact. 7, CSCW1 (April 2023), 1–38.
- Wadinambiarachchi et al. (2024) Samangi Wadinambiarachchi, Ryan M Kelly, Saumya Pareek, Qiushi Zhou, and Eduardo Velloso. 2024. The effects of generative AI on design fixation and divergent thinking. In Proceedings of the CHI Conference on Human Factors in Computing Systems. ACM, New York, NY, USA, 1–18.
- Wan et al. (2024) Ruyuan Wan, Simret Gebreegziabhe, Toby Jia-Jun Li, and Karla Badillo-Urquiola. 2024. CoCo Matrix: Taxonomy of cognitive contributions in co-writing with intelligent agents. arXiv [cs.HC] (May 2024).
- Wang et al. (2020) Dakuo Wang, Elizabeth Churchill, Pattie Maes, Xiangmin Fan, Ben Shneiderman, Yuanchun Shi, and Qianying Wang. 2020. From human-human collaboration to human-AI collaboration. In Extended Abstracts of the 2020 CHI Conference on Human Factors in Computing Systems. ACM, New York, NY, USA.
- Weld and Bansal (2019) Daniel S Weld and Gagan Bansal. 2019. The challenge of crafting intelligible intelligence. Commun. ACM 62, 6 (May 2019), 70–79.
- Wiethof et al. (2021) Christina Wiethof, Navid Tavanapour, and Eva Bittner. 2021. Implementing an intelligent collaborative agent as teammate in collaborative writing: Toward a synergy of humans and AI. In Proceedings of the 54th Hawaii International Conference on System Sciences. Hawaii International Conference on System Sciences.
- Wilbers et al. (2024) Simon Wilbers, Johanna Gröpler, Bastian Prell, and Jörg Reiff-Stephan. 2024. Overall writing effectiveness: Exploring students’ use of LLMs, pushing the limits of automated text generation. In Smart Technologies for a Sustainable Future. Springer Nature Switzerland, Cham, 11–22.
- Yang et al. (2022) Daijin Yang, Yanpeng Zhou, Zhiyuan Zhang, Toby Jia-Jun Li, and L C Ray. 2022. AI as an active writer: Interaction strategies with generated text in human-AI collaborative fiction writing 56-65. (2022), 56–65.
- Yang et al. (2019) Qian Yang, Justin Cranshaw, Saleema Amershi, Shamsi T Iqbal, and Jaime Teevan. 2019. Sketching NLP: A case study of exploring the right things to design with language intelligence. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems. ACM, New York, NY, USA.
- Zhang et al. (2020) Yunfeng Zhang, Q Vera Liao, and Rachel K E Bellamy. 2020. Effect of confidence and explanation on accuracy and trust calibration in AI-assisted decision making. arXiv [cs.AI] (Jan. 2020).
- Zhang et al. (2023) Zheng Zhang, Jie Gao, Ranjodh Singh Dhaliwal, and Toby Jia-Jun Li. 2023. VISAR: A human-AI argumentative writing assistant with visual programming and rapid draft prototyping. In Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology. ACM, New York, NY, USA, 1–30.
Appendix A Appendix: Systematic Review Screening Questions
We provide the inclusion/exclusion used for the first screening in Table 4, and the QA Checklist used for the second screening in Table 5.
| Type | Criteria |
| Inclusion | The paper needs to present or build on original social science theories or framework related to collaborative writing or dialogue |
| The paper is written in English | |
| The paper focuses on English writing or English dialogue | |
| Exclusion | Not related to collaborative writing |
| The paper focuses on the pedagogical side of collaborative writing in education or collaborative learning | |
| The paper focuses on cross-language collaborative writing, such as second language learners, writing in non-English languages, translation or interpretation | |
| The paper focuses on story or fictional writing | |
| The paper focuses on authorship or publishing ethics | |
| The paper focuses on techniques for version control of documents | |
| Proceedings titles | |
| Unavailable for retrieval | |
| Unpublished (i.e., “Early Access” papers on WoS) | |
| Duplicated |
| Question | Scoring |
| Does the paper introduce or build on original theories or frameworks related to collaborative writing? | • Yes (=1 point) • Partially (=0.5 point) • No (=0 point) |

### v3

@@
 Abstract.
-In this paper, we conduct a critical review of existing theories and frameworks on human-human collaborative writing to assess their relevance to the current human-AI paradigm in professional contexts, and draw seven insights along with design implications for human-AI collaborative writing tools.
-We found that, as LLMs nudge the writing process more towards an empirical “trial and error” process analogous to prototyping, the non-linear cognitive process of writing will stay the same, but more rigor will be required for revision methodologies.
-This shift would shed further light on the importance of coherence support, but the large language model (LLM)’s unprecedented semantic capabilities can bring novel approaches to this ongoing challenge.
-We argue that teamwork-related factors such as group awareness, consensus building and authorship – which have been central in human-human collaborative writing studies – should not apply to the human-AI paradigm due to excessive anthropomorphism.
-With the LLM’s text generation capabilities becoming essentially indistinguishable from human-written ones, we are entering an era where, for the first time in the history of computing, we are engaging in collaborative writing with AIs at workplaces on a daily basis.
+In this paper, we conduct a critical review of existing theories and frameworks on human-human collaborative writing to assess their relevance to the current human-AI paradigm in organizational workplace settings, and draw seven insights along with design implications for human-AI collaborative writing tools.
+Our main finding was that, as we delegate more writing to AI, our cognitive process shifts from the traditional planning/translating/reviewing process to a planning/waiting/reviewing process, breaking the process due to the waiting that occurs in between.
+To ensure that our cognitive process remains intact, we suggest a “prototyping” approach, where the tool allows for faster iterations of the cognitive process by starting with smaller chunks of text, and gradually moving on to a fully fleshed-out document.
 We aim to bring theoretical grounding and practical design guidance to the interaction designs of human-AI collaborative writing, with the goal of enhancing future human-AI writing software.
 1. Introduction
-As large language models (LLMs) increasingly become our daily co-writers in a uniquely asymmetrical relationship unlike any we have previously encountered, there is a need to reassess foundational theories and frameworks of collaborative writing to adapt to the human-AI paradigm.
-Their capabilities in natural language generation are becoming indistinguishable from human writing (Hämäläinen et al., 2023; Clark et al., 2021), but the mechanism of human-AI collaborative writing is clearly distinct from a human-human one.
-To reflect this paradigm shift, new frameworks and taxonomies are emerging, from domain-specific taxonomy development (Lee et al., 2024b), recognizing human-side effort through entropy and information gain (Wan et al., 2024), getting inspiration from the Overall Equipment Effectiveness Model (Wilbers et al., 2024), to exploring a non-linear framework to go beyond the linear human-only/human-AI/AI-only dimension of collaboration modes.
+When humans made the shift from writing on paper to typing on word-processors in the 1980s, studies found that writers engaged in less conceptual planning, reread texts more extensively, and there was poorer recall of their own text (Haas, 2013).
+How then, will writing with Large Language Models (LLMs) affect the way that we write?
+As LLMs increasingly become our daily co-writers, and as their capabilities in natural language generation become increasingly indistinguishable from human writing (Hämäläinen et al., 2023; Clark et al., 2021; Zhu et al., 2024), we see a need to reassess foundational theories and frameworks on collaborative writing, and see how they adapt to the human-AI paradigm.
+To reflect this paradigm shift, new frameworks and taxonomies of collaborative writing are emerging, from domain-specific taxonomy development (Lee et al., 2024b), recognizing human-side effort through entropy and information gain (Wan et al., 2024), getting inspiration from the Overall Equipment Effectiveness Model (Wilbers et al., 2024), to exploring a non-linear framework to go beyond the linear human-only/human-AI/AI-only dimension of collaboration modes.
 (Hutson, 2025).
-While many of these proposals build on a partial aspect of existing theories and frameworks of collaborative writing, we believe that this must be approached holistically, synthesizing what past studies have taught us about the mechanisms of collaborative writing.
-It goes without saying that legacy theories and frameworks on collaborative writing have (naturally) focused only on the human-human mode.
-While the already established understandings of human-human collaborative writing should certainly apply to human-AI to some degree, we believe now is the moment to reassess our understanding of collaborative writing so far, and adapt it to the ever increasing human-AI mode.
+While many of these proposals build on a partial aspect of existing theories and frameworks on collaborative writing, we believe that this must be approached holistically, synthesizing what past studies have taught us about the mechanisms of collaborative writing.
+Naturally, legacy theories and frameworks on collaborative writing have all focused on humans writing with humans.
+To what degree do these understandings of human-human collaborative writing apply to the human-AI paradigm?
+We believe now is the moment to look back to the diverse literature in collaborative writing, and see what we can take from them to the increasing human-AI mode.
 Hence our research question: what insights can we draw from theories/frameworks on human-human collaborative writing to apply to the design of human-AI writing tools?
@@
 In this paper we conduct a critical review of legacy theories and frameworks on collaborative writing, and analyze what we see to be relevant to the human-AI paradigm as well as what only applies to the human-human paradigm, bringing clarity to the theoretical groundings required for the design of future human-AI collaborative writing tools.
-Our main findings are that, given the shift in the human’s role from hands-on writer to an editorial decision maker, the non-linear cognitive process of writing will stay the same, but more attention will be required for rigorous revision methodologies.
-In turn, this may reduce the focus on the planning aspects of collaborative writing, nudging writing more towards an empirical “trial and error” process than a fully planned out one, analogous to prototyping.
-This shift would require more focus on supporting coherence between human and AI writing, but we believe that the LLM’s unprecedented semantic capabilities can bring a break through to this ongoing challenge.
-Teamwork-related factors such as group awareness, consensus building and authorship, which have been central in human-human collaborative writing studies, will not apply to the human-AI paradigm due to excessive anthropomorphism, but it is equally the AI co-writer’s role to facilitate those factors between human writers to offload any unnecessary cognitive load.
+Our main findings are that, when the human’s role shifts from a hands-on writer to an editorial decision maker, the cognitive process of writing (Flower and Hayes, 1981) shifts from the traditional planning/translating/reviewing process to planning/waiting/reviewing (Figure 1), making the cognitive process discontinuous.
+For this we suggest AI tools to streamline the cognitive process by shortening the waiting stage as well as accelerating the overall cycle through a “prototyping” approach to writing, such as enabling human writers to revise AI generated text faster, and directly connecting that as the next set of prompts to test new ideas.
+Coherence support between human and AI writing is another ongoing challenge, but the LLM’s strong semantic capabilities can provide real-time analyses that were previously difficult with traditional methods.
+Teamwork-related factors such as group awareness, consensus building and authorship, which have been central in human-human collaborative writing studies, will not apply to the human-AI paradigm due to excessive anthropomorphism, but it is the AI co-writer’s role to facilitate such factors for human writers to offload cognitive load.
 Finally, be it human-written or AI-written, writing will remain to be socially constructed by readers, which suggests that the consideration of reader profiles is inseparable from the text generation process.
 We provide practical design implications and example interface sketches with our insights to provide concrete design directions.
+The main contributions of this work are as follows:
+We define the limitations of existing human-human collaborative writing theories and frameworks when applied to the human-AI paradigm
+We propose an updated cognitive model when collaborative writing with AI and discuss its implications
+We provide actionable design implications for future human-AI writing tools based on what we can apply from collaborative writing theories
 1.1. Study scope
 The scope of this paper is limited to the pragmatic support of writing tasks in organizational workplace settings.
-This includes cases such as clinicians drafting discharge summaries, course coordinators drafting a response to student questions on online forums, meeting participants drafting a summary to send out to their team, or academic staff crafting grant proposals.
-We do not consider in fictional writing or rhetoric, nor do we consider topics around authorship, emotional satisfaction, or the deskilling aspects of writing.
-In particular, collaborative writing has traditionally had much focus on ownership and authorship, but building on Schneiderman (Shneiderman, 2022) we take the stance that LLMs are and should merely be a tool to help us get the job done in the most effective manner, and avoid any unnecessary forms of anthropomorphism in framing the AI as a “team mate”.
+A prototypical example of such case would be drafting emails with the help of LLMs.
+Other cases include clinicians drafting discharge summaries, course coordinators drafting a response to student questions on online forums, meeting participants drafting a summary to send out to their team, or academic staff crafting grant proposals.
+We do not consider fictional writing or rhetoric (Yang et al., 2022; Calderwood et al., 2020), nor do we consider the deskilling aspects of writing (Kosmyna et al., 2025).
+In particular, studies on collaborative writing have traditionally focused much on ownership and authorship(Ede and Lunsford, 2001; Lunsford and Ede, 1986; Stokel-Walker, 2023), but building on the views of Naikar et al.
+(Naikar et al., 2025) and Schneiderman (Shneiderman, 2022) we take the stance that LLMs are and should merely be a tool to help us get the job done in the most effective manner, and avoid any unnecessary forms of anthropomorphism that frames the AI as a “team mate”.
 2. Background
@@
 2.1. Social Computing Paradigm
-The Computers As Social Actors (CASA) paradigm (Nass et al., 1994), which revealed that humans apply the same social rules they would apply to humans to computers, has been a dominant framework in Human Computer Interaction (HCI).
+The Computers As Social Actors (CASA) paradigm (Nass et al., 1994), which revealed that humans often apply social rules to computers, has been a dominant framework in Human Computer Interaction (HCI).
 The underlining premise is simple: the way humans interact with computers is inherently social.
-The social response theory (Nass and Moon, 2000) took this a step further, explaining that such social rules are “mindlessly applied”, and that alternative explanations such as anthropomorphism simply cannot explain this phenomenon.
+The social response theory (Nass and Moon, 2000) took this a step further, explaining that such social rules are often “mindlessly applied” (Nass and Moon, 2000), and that alternative concepts such as anthropomorphism are not sufficient to explain this phenomenon.
 Hence, to design interactions for new technologies, it is best to reframe computer interaction as a social interaction, and appropriately lean on relevant studies in social sciences along the way.
@@
 For example, in the field of Explainable AI, Miller (Miller, 2019) suggests that we build on what the social sciences have already uncovered about human-to-human explanation mechanisms, because after all, “explanations are social”, and being in a vacuum of computer science literature would not solve the problem.
-Recently, such an approach has begun to become evident in dialogue studies, as represented in the emerging field of Human-Machine Communication (HMC), an area of communication research that regards technology as “communicative subjects, instead of mere interactive objects” (Guzman and Lewis, 2020).
+Recently, such an approach has become evident in dialogue studies, as represented in the emerging field of Human-Machine Communication (HMC), an area of communication research that regards technology as “communicative subjects, instead of mere interactive objects” (Guzman and Lewis, 2020).
 That is, HMC treats machines no longer as mediators, but as actors directly exchanging messages with humans.
-This reframing expands the focus of communication studies from human-human only to include human-AI contexts.
-For example, while it is well known that the lexical, syntactic and semantic alignment between two people engaged in a dialogue is crucial for mutual understanding (Pickering and Garrod, 2004), Ostrand and Berger claim that such alignment should be properly considered when building LLMs as well (Ostrand and Berger, 2024) for improved usability and personalization.
-This has also led to interesting discussions on the effects such alignment would have on bias amplification and echo chambers (Knoeferle et al., 2025).
-With collaborative writing, however, there is still little work at the intersection of traditional human-human collaborative writing and the emerging human-AI writing paradigm.
-This is perhaps due to the fact that, until recently, writing with AI primarily referred to assistance tools such as spell- or grammar-checking.
-It also remains unclear as to what extent writing with LLMs can be considered “collaborative writing”.
-However, when a human tasks an LLM to write up a draft of a document, the human-AI pair is working “toward coordinated consensus that is reflected in a document that is written by one of the team members”, which we see to be undeniably qualifying as a form of collaborative writing (Lowry et al., 2004) (it is also worth noting that, according to Posner and Baecker (Posner and Baecker, 1992), single writer strategy was actually the most common strategy adopted even in human-only groups).
-If that is the case, the underlining reframing that is needed here remains the same.
-Writing is hardly a solitary act, be it with humans or with AI: it happens through the participation and interaction with others (Ede and Lunsford, 2006).
-We can see here that writing, too, is inherently social.
-As Wang et al.
-(Wang et al., 2020) suggest, it is crucial to integrate existing theories from HCI and Computer-Supported Cooperative Work (CSCW) in designing for the optimal form of human-AI collaborative writing.
+This reframing expands the focus of communication studies beyond human-human contexts to human-AI contexts.
+For example, while it is well known that the lexical, syntactic and semantic alignment between two people engaged in a dialogue is crucial for mutual understanding (Pickering and Garrod, 2004), Ostrand and Berger claim that such alignment should be considered when building LLMs as well (Ostrand and Berger, 2024), for improved usability and personalization.
+This has also led to interesting discussions on the effects that such alignment would have on bias amplification and echo chambers (Knoeferle et al., 2025).
+With collaborative writing, however, there is still little work that applies what we know from human-human collaborative writing to the emerging human-AI writing paradigm.
+This is partially due to the fact that, until recently, writing with AI primarily referred to limited assistance tools such as spell- or grammar-checking.
+LLMs have brought a turning point to this field with their impressive text generation and semantic analysis capabilities.
+Writing tasks in educational and communication contexts are now the second most frequent usage of Claude (Handa et al., 2025) (the first being coding).
+Inspired by the CASA paradigm, the way humans write with AI could also be inherently social, which leads us to this systematic review to see which aspects of existing studies on human-human writing will apply to human-AI writing.
 2.2. Collaborative Writing
-In 1969, Michel Foucault questioned the notion of the solitary author in the book “What is an author?”, stating that the author’s role is to regulate meaning and the circulation of text in an ever changing society (Foucault, 1979).
-The cognitive process theory of writing (Flower and Hayes, 1981), a widely adopted theory for unveiling an individual’s writing process, was updated by Hayes 15 years later (Hayes, 1996) to add more emphasis to the “social” aspect of writing, incorporating readers and collaborators into their framework.
+While writing was once thought to be a solitary act (Foucault, 1979), there was a gradual realization that, in many cases, humans in fact write together, leading to the field of collaborative writing.
+To reflect this, Flower and Hayes’ widely adopted theory of an individual’s cognitive writing process (Flower and Hayes, 1981) was updated by Hayes 15 years later (Hayes, 1996), emphasizing to the “social” aspect of writing and incorporating collaborators as well as readers into their framework.
 Ede and Lunsford (Lunsford and Ede, 1986) were the among the first to initiate collaborative writing as an academic research topic in 1986, reporting that 87 percent of respondents from six major professions sometimes wrote as a team or a group.
@@
 (Lowry et al., 2004) established the taxonomy of collaborative writing, outlining not just the various roles but also the writing strategies that groups adopt when writing together.
-As the role of computers and software becomes increasingly important in writing, this taxonomy is still getting updated to this day, including Collage (Buschek, 2024), CoCo matrix (Wan et al., 2024), Overall Writing Effectiveness (Wilbers et al., 2024), neurodiversity-inspired non-linear framework (Hutson, 2025), taxonomy of “text functions” (Larsen-Ledet et al., 2020) to name a few.
+As the role of computers and software are becoming increasingly important in writing, this taxonomy is getting updated to this day, including Collage (Buschek, 2024), CoCo matrix (Wan et al., 2024), Overall Writing Effectiveness (Wilbers et al., 2024), neurodiversity-inspired non-linear framework (Hutson, 2025), taxonomy of “text functions” (Larsen-Ledet et al., 2020) to name a few.
 Our work builds on these studies of emerging taxonomies, clarifying the limitations of existing human-human taxonomies and suggesting what needs to be renewed.
-LLMs have brought a turning point to this field with their impressive text generation and semantic analysis capabilities.
-Given that writing tasks in educational and communication contexts are the second most frequent usage of Claude (Handa et al., 2025), an unprecedented number of workers are engaging in some form of collaborative writing with AIs on a daily basis.
-We believe it is more important now than ever to unveil the mechanisms of collaborative writing, especially between humans and AI.
+It also remains unclear as to what extent writing with LLMs can be considered “collaborative writing”.
+However, when a human tasks an LLM to write up a draft of a document, the human-AI pair is working “toward coordinated consensus that is reflected in a document that is written by one of the team members” (Lowry et al., 2004), which qualifies as a form of collaborative writing (it is also worth noting that, according to Posner and Baecker (Posner and Baecker, 1992), single writer strategy was actually the most common strategy adopted, even in human-only groups).
+If that is the case, it seems viable to consider writing with AI to be a legitimate form of collaborative writing.
+Be it with humans or with AI, writing is hardly a solitary act: it happens through the participation and interaction with others (Ede and Lunsford, 2006).
 2.3. Human-AI collaboration
-Achieving synergy between human and AI performance has been the core focus of human-centered AI and human-robot interaction studies.
-One example is Explainable AI (XAI), attempting to achieve complementary performance between humans and AI for various tasks such as decision making and classification, through the transparency of the otherwise black box AI.
-Allowing users to drill down and ask follow up questions until they are satisfied has been an often discussed yet technically difficult challenge in XAI (Weld and Bansal, 2019), but LLMs have dramatically enabled us to study this approach further.
-In general, however, the various approaches in XAI have led to mixed results with regards to its effectiveness on performance (Miller, 2023), which leads to Vaccaro et al.’s (Vaccaro et al., 2024) impressive meta-analysis suggesting that creation-related tasks (including writing) with GenAI statistically have more potential for achieving synergy– the positive complementary outcome from human-AI collaboration – and that the appropriate allocation of tasks between humans and AI may be the key to further progression.
+Achieving synergy between human and AI performance has been a primary focus of human-centered AI and human-robot interaction studies.
+One example is Explainable AI (XAI), which, among other goals, seeks to achieve complementary performance between humans and AI in tasks such as decision making and classification by increasing the transparency of otherwise black-box models.
+Allowing users to drill down and ask follow up questions until they are satisfied has been an often discussed yet technically difficult challenge in XAI (Weld and Bansal, 2019), but LLMs have dramatically enabled us to study this approach further (Ma et al., 2024).
+In general, however, the various approaches in XAI have led to mixed results with regards to its effectiveness on performance (Miller, 2023), which leads to Vaccaro et al.’s (Vaccaro et al., 2024) impressive meta-analysis suggesting that creation-related tasks with GenAI (including writing) have more statistical potential for achieving synergy – the positive complementary outcome from human-AI collaboration – and that the appropriate allocation of tasks between humans and AI may be the key to further progression.
 This is in line with past work (Zhang et al., 2020) which found that showing AI confidence or local feature explanations had no or detrimental effect to decision accuracy, suggesting that XAI can be most effective when the error boundaries of human’s and AI’s are as distinct as possible.
@@
 Similar to Yang et al.
-(Yang et al., 2022), we intentionally separate AI-assisted writing from AI as the co-writer.
+(Yang et al., 2022), we intentionally distinguish AI-assisted writing from AI as the co-writer.
 Existing studies for AI as a co-writer have mostly focused on creative writing (Clark et al., 2018; Yang et al., 2019; Calderwood et al., 2020; Yang et al., 2022).
 Clark et al.
-(Clark et al., 2018) suggests that by reducing the intrusiveness of AI assistance for creative writing and adopting a pull method for the interactions, it can “allow people to write more closely to their normal writing process.” Yang et al.
-(Yang et al., 2022) took a turn-taking approach in building a “finish the story” system, finding that when users are equipped with the ability to edit AI generated text, they require only partial readability for the AI outputs because they can go in and fix the fluency of the overall text themselves.
+(Clark et al., 2018) suggest that by reducing the intrusiveness of AI assistance for creative writing and adopting a pull method for interactions, it can “allow people to write more closely to their normal writing process.” Yang et al.
+(Yang et al., 2022) took a turn-taking approach in building a “finish the story” system, finding that when users are equipped with the ability to edit AI generated text, they require only partial readability from AI because they can go in and fix the fluency themselves.
 At the same time, in the adjacent field of image generation, Wadinambiarachchi et al.
@@
 (Lee et al., 2022) proposes CoAuthor, a large dataset of recorded interactions between writers and GPT-3 across 1445 writing sessions, which has paved the way for HCI researchers to analyze human-AI writing cases in detail.
-Our review builds on this fundamental Human-AI collaboration research to draw out the best of human and AI performance, specifically for writing tasks in organizational workplace settings, and in a way that is rooted in existing theories and frameworks.
+Our review builds on these fundamental Human-AI collaboration research, and seek to understand how human-AI writing in organizational workplace settings can be designed to draw out the best of human and AI performance.
 To the best of our knowledge, this is the first work to holistically reassess legacy collaborative writing theories and frameworks for the LLM era.
@@
 (Moher et al., 2009), we selected 2 databases (Web of Science and the ACM Digital Library) and used the queries in Table 1.
-Because the selection of the ACM Digital Library was primarily to address RQ2, we added an additional filtering for LLM related papers, which was not necessary for Web of Science due to its broader, non-HCI-specific scope.
-Given that collaborative writing has been studied for decades, we did not limit by publication year.
+It is important to note that this study aims to bridge existing collaborative writing theories/frameworks (which fall under various research fields, such as rhetoric and writing, cognitive science, and Computer-Supported Cooperative Work) with AI writing studies (which are often from HCI).
+For this reason, we relied on the single phrase “collaborative writing” for the Web of Science query to cast an intentionally wide net - we wanted to gather as much existing work from outside the fields of HCI and computer science, new and old.
+At the same time, we intentionally added a filtering for LLM related keywords for the ACM library query, because the selection of the ACM library was primarily to address RQ2.
+This was not necessary for Web of Science due to its broader, non-HCI-specific scope.
+Given that collaborative writing has been studied for decades, we also did not limit by publication year.
 | Database | Query |
@@
 After removing 7 duplicates, we applied the inclusion/exclusion criteria in Appendix A to the 3,060 entries based on the titles and abstracts.
+In addition to fictional writing, rhetoric, and authorship, we also excluded studies in the educational context, such as those related to the pedagogy of collaborative writing or second language learners.
+This is because, in the context of organizational workplace settings, we see such considerations to be largely irrelevant – the primary aim for such users is to simply get the task done with less time and effort.
+As mentioned in Section 1.1, the scope of this paper is limited to the pragmatic support of writing tasks in organizational workplace settings, such as writing emails, clinicians drafting discharge summaries, meeting participants drafting a summary to send out to their team, or academics crafting grant proposals.
+While educational approaches in these settings would certainly be of benefit in theory, in most contexts, it would get in the way of the users’ objectives.
 This left us with 65 entries – we lightly reviewed the full text, and screened with the predefined QA checklist in Appendix A.
@@
 | Zhang et al. (Zhang et al., 2023) | 2023 | Builds on Toulmin’s argument model and the Cognitive Process Theory of Writing |
-To ensure an objective process, two reviewers (the first and second author) took the same 15 randomly selected entries and applied the inclusion/exclusion criteria screening to check agreement levels.
+To reduce potential bias, two reviewers (the first and second author) took the same 15 randomly selected entries and applied the inclusion/exclusion criteria screening to check agreement levels.
 This sample screening not only revealed substantial agreement between the 2 reviewers (Cohen’s ), it also helped to add specificity to the exclusion criteria, which the first author applied to the rest of the entries.
@@
 As obvious as this may sound, LLMs need no rest.
-They are “always” available (except due to technical downtime, etc.), and there is no need for social hesitation or awkwardness when asking for help.
-As Gero et al.
-(Gero et al., 2023)’s work revealed, writers feel there is no need to work hard to establish or maintain a good relationship with AI agents, which frees them from the worry of being judged.
-While this may seem contrary to the CASA paradigm, Ede and Lunsford (Ede and Lunsford, 2006) interestingly point out that in the most ideal form of collaborative writing, you should “let everything show - the stupidities, the backsliding and illogicalities, and everything else that goes into writing a text - as well as to be open to those things in the other person, and to the other person’s strengths”.
-This suggests that, if used appropriately, the lack of social norm could paradoxically be what allows for exceptionally effective collaboration, be it for human-human or human-AI.
-In other words, writers are able to ask from LLMs much more than what they would request from humans, without worrying about social norms.
+They are software, and are “always” available (except due to technical downtime, etc.).
+Additionally, as Gero et al.
+(Gero et al., 2023)’s work revealed, writers feel there is no need to establish or maintain a good relationship with AI agents, which frees them from the worry of being judged.
+As a result, writers are able to ask from LLMs much more than what they would request from humans, without worrying about social norms.
 For example, Laban et al.
-(Laban et al., 2024) experimented with executable edits which allowed authors to edit faster than simple chat interactions.
-There is much to borrow from AI-powered code editors such as Cursor222See https://www.cursor.com or Github Copilot333See https://github.com/features/copilot in this aspect, where it is common for suggestions to be applied to the codebase with a click, beyond suggestions via chat.
+(Laban et al., 2024) experimented with executable edits which allowed authors to go beyond typical chat interactions and edit faster.
+There is much to learn from AI-powered code editors such as Cursor222See https://www.cursor.com or Github Copilot333See https://github.com/features/copilot in this regard, where it is already common for suggestions to be applied to the codebase with a click, beyond suggestions via chat.
 4.1.2. No need for consensus building
@@
 Since AI tools have no opinions, values, or goals, we believe there is no need to consider building consensus with AI.
-However, there is value in human writers making their values and goals clear to an AI writing tool, and for human co-authors to do so with potential support from an AI writing tool, as we will discuss in our findings.
+However, there is still value in human writers communicating their values and goals clearly to an AI writing tool, and for a team of human co-authors to share among themselves with support from an AI writing tool, as we will discuss in our findings.
 4.1.3. Distinguish content from instructions
@@
 4.1.4. Distrust is encouraged
-As meaningless as it may be to state, machines don’t care.
-Therefore, a certain level of distrust and skepticism towards AI as a co-writer is, ironically, encouraged.
-This is clearly different from human-human collaborative writing, where the establishment of trust and camaraderie is seen as a major component for success (Ens et al., [n.
+As meaningless as it may be to state, machines do not care.
+Therefore, a certain level of distrust and skepticism towards AI as a co-writer is encouraged.
+Jacovi et al.
+(Jacovi et al., 2021) distinguish trust with trustworthiness, stating that distrust between human and AI could in fact be desirable, in conditions where the humans are aware of the AI’s non-trustworthiness.
+This seems to be different from human-human collaborative writing literature, where the establishment of trust and camaraderie is seen as a major component for success (Ens et al., [n.
 d.]; Lowry et al., 2004).
-We build on the approach of Cai et al.
+We build on the approach that Cai et al.
 (Cai et al., 2019a) and Gu et al.
-(Gu et al., 2021) in creating clinical decision support tools: instead of waiting for AI to be good enough to be “fully automatable one day” (Gu et al., 2021), we should accept its “imperfection” and move towards creating tools with careful HCI considerations, which, in turn, would lead to better calibrated trust (Jacovi et al., 2021) and utility while experiencing less workload (Cai et al., 2019a).
+(Gu et al., 2021) took in creating AI decision support tools: instead of waiting for AI to be good enough to be “fully automatable one day” (Gu et al., 2021), we should accept its “imperfection” and move towards creating tools with careful HCI considerations, which, in turn, would lead to better calibrated trust (Jacovi et al., 2021) and utility while experiencing less workload (Cai et al., 2019a).
 4.2. Seven insights and design implications from legacy theories/frameworks
 We now introduce the seven insights we drew from our review, along with design implications for future human-AI collaborative writing tools. The results are presented in Table 3.
-| Theme | Design implication |
-| Writing process and strategies are iterative, non-linear and interchangeable any time | Allow for edits or additional prompts at any given moment during the process of writing |
-| Supporting planning has been the primary focus in collaborative writing tools | Equip users with granular revision control |
-| Design novel methods to effectively prototype complex documents | |
+| Themes from human-human collaborative writing | Design implications for human-AI collaborative writing |
+| Writing processes and strategies are iterative, non-linear and interchangeable any time | Shift to a “prototyping” approach of text to streamline the discontinuous planning/waiting/reviewing process |
+| Supporting planning has been the primary focus in collaborative writing tools | More rigor is required for helping human writers review and revise AI generated text |
 | Group awareness is a key to success | Provide a moment for the writer to reflect on their expectations and needs from the AI co-writer |
@@
 | Lack of clarity with authorship | Design data structures and visual cues to distinguish human-written text from AI-written ones |
-| Provide built-in verification methods for AI-generated text | |
 | Collaborative writing tools serve as external cognition | Make sure that the AI co-writer is contributing to both the production of text as well as the facilitation of the process itself |
-4.2.1. Writing process and strategies are iterative, non-linear and interchangeable any time
+4.2.1. Writing processes and strategies are iterative, non-linear and interchangeable any time
 Cognitively, writing processes are highly flexible and non-linear.
 This is one of the key insights of the Cognitive Process Theory of Writing (Flower and Hayes, 1981), which was the first to model the cognitive process of writing and has been referred to by many collaborative writing studies since.
-They proposed a hierarchically-organized model that primarily consists of 3 steps: (1) planning; (2) translating; and (3) reviewing, but writers “do not march through these processes in a simple 1, 2, 3 order” (Flower and Hayes, 1981).
-The writing process is essentially a constant iteration of creating and revising goals and sub-goals, and therefore key subprocesses such as idea generation, revision and evaluation can and should interrupt other processes at any given moment during the process of writing.
-More simply put: revision is not something that happens after writers do their writing.
-Rather, it happens “while in the midst of generating a text” (Faigley and Witte, 1981).
-This non-linear and interchangeable nature of the writing process is a common theme across various theories and frameworks.
-Ede and Lunsford (Ede and Lunsford, 2006), when explaining their proposed model of hierarchical and dialogic modes of writing, equally stress the importance of resisting the urge to treat writing as a simple dichotomy between the two: “it is inherently mixed and paradoxical”.
+They proposed a hierarchically-organized model that primarily consists of 3 steps: (1) planning; (2) translating; and (3) reviewing, with the monitoring process functioning as an orchestrator that decides when the writer moves from one process to another (Figure 1).
+A key aspect of this theory is that writers “do not march through these processes in a simple 1, 2, 3 order” (Flower and Hayes, 1981) - rather, it is a constant iteration of creating goals, writing towards that goal, reading what was written, revising goals, writing towards the new goal, and so forth.
+Therefore, key subprocesses such as idea generation, revision and evaluation are interruptable by other processes at any given moment during the writing process.
+More simply put: writers revise as they write, plan as they revise, and write as they plan.
+Contrary to many beliefs, revision is not something that happens after writers do their writing, but rather something that happens “while in the midst of generating a text” (Faigley and Witte, 1981).
+This non-linear and interchangeable nature of the writing process was a common theme across various theories and frameworks.
+When explaining their proposed model of hierarchical and dialogic modes of writing, Ede and Lunsford (Ede and Lunsford, 2006) equally stress the importance of resisting the urge to treat writing as a simple dichotomy between the two: “it is inherently mixed and paradoxical”.
 Even prior to Google Docs, collaborative writing tools have supported parallel-partitioned writing, allowing members to view and edit each other’s work at any given moment, which is known to increase teamwork effectiveness (Lowry and Nunamaker, 2003).
-Design implication #1: Allow for edits or additional prompts at any given moment during the process of writing
-Even if AI co-writers become the ones that generate most of the actual text, it is still vital for collaborative writing tools to respect and support the human’s cognitive processes.
-While we found studies that quite literally build on the 3 steps of the Cognitive Process Theory of Writing and incorporate its non-linearity (Chakrabarty et al., 2024; Zhang et al., 2023), we feel that a key design implication here is the importance of allowing the user to seamlessly revise and test new ideas at the same speed as when they are writing themselves.
-That is, the user should not be forced to wait for the LLM’s text generation to complete.
-Rather, the generation should always run in the background and be interruptable, so that the user can make edits or request for additional prompts at any given moment during the collaborative writing process.
-For instance, imagine an interface where the text generation is constantly run in the background, with no explicit “generate” button.
-We see this to be much more in line with the human’s non-linear and interchangeable cognitive process of writing, compared to a chat based interaction where the reviews, edits and imaginations are forced to take place in a linear manner.
+Design implication #1: Shift to a “prototyping” approach of text to streamline the discontinuous planning/waiting/reviewing process
+Even if AI co-writers become the ones that generate most of the actual text, it is natural to assume the same cognitive processes will be at play for the human writer, and this should be respected and supported by AI writing tools.
+However, the non-linear, interchangeable nature of the cognitive process of writing is dependent on the fact that the writers are writing the text by themselves, or at least are fully aware of the written content.
+In the cases where the AI generates a majority of (if not the entire) the writing, the traditional planning/translating/reviewing process essentially shifts to a planning/waiting/reviewing process, thereby making the human writer’s cognitive process discontinuous due to the waiting that occurs in between.
+In such cases, we believe it is crucial that the AI writing tool acts as an active enabler of the human’s cognitive process.
+While we found studies that build on the three stages of the Cognitive Process Theory of Writing with AI (Chakrabarty et al., 2024; Zhang et al., 2023), we argue that such literal implementations act as a hindrance to the iterative interchangeable nature of the human cognitive process.
+Rather, AI tools should work to (1) shorten the waiting process as much as possible and (2) accelerate the cognitive process, to allow human writers to test new ideas at a similar cognitive speed as when they are writing themselves.
+This is where the analogy of prototyping comes in.
+In a prototyping process, the speed at which you iterate is key: you build something rough, test quickly, learn what could be improved, and move on to the next cycle.
+We argue that AI writing tools would be more effective if they support the “prototyping” of text.
+For instance, imagine an interface where the text generation is intentionally provided as bite-size, digestible chunks which the human writer can read and review with little time and effort, compared to a full size paragraph.
+The human writer can directly reorder those chunks, add additional text or instructions for further refinements and run the generation again to retrieve another round of digestible chunks, thereby ”prototyping” the writing until they are satisfied and hit the “develop full document button” for the final version.
+Buschek’s work on comparing the current human-AI collaborative writing scene with literary collage from avant-garde literature (Buschek, 2024) frames this approach eloquently: writing with LLMs is very much the assembling of “prefabricated text”.
+Such prototyping approaches can potentially lead to the mitigation of the loss of critical thinking (Lee et al., 2025) in human writers, which are already seen in essay writing tasks (Kosmyna et al., 2025) or with image generation tools (Wadinambiarachchi et al., 2024).
+If we are to blindly generate documents or emails to send to our peers hardly reviewing their output, the results would be catastrophic, both for our cognitive process, and for workplace collegiality and responsibility.
+We argue that the problem here is not the ease of text generation, but the difficulty in reviewing and revising fully developed text.
+We see this to be a highly interesting and novel challenge that we welcome researchers to build further.
 4.2.2. Supporting planning has been the primary focus in collaborative writing tools
 Traditionally, supporting planning has been the primary focus in CW tools (De Silva, 2007; Zhang et al., 2023; Lowry and Nunamaker, 2003).
-Prasetyo and Bandung discuss the planning features of their e-book writing tool, but features for reviewing hardly go beyond simple annotations and comments.
-This is understandable, since the translation phase is mostly the humans doing their work, and revision requires the humans to read each other’s work: they were both primarily human-centric tasks.
-However, particularly in cases where the LLMs take the role of writer, we question the necessity of such emphasis on planning.
+Prasetyo and Bandung discuss the planning features of their e-book writing tool in depth, but features for reviewing hardly go beyond simple annotations and comments.
+This is understandable, since revision was primarily a human-centric task: it requires humans to read the work that they wrote.
+However, in cases where the LLMs take the role of writer, we question the necessity of such emphasis on planning as a function of the tool.
 Indeed, in Chakrabarty et al.’s study on supporting professional writers with LLMs (Chakrabarty et al., 2024), the AI’s contributions with the lowest retention rate (i.e.
@@
 The strength of the AI co-writer lies not in the meticulous refinement of the message to convey, but simply in the faster and easier translation of ideas to words, as well as rigorous revision of formatting and errors.
-Interestingly, this shift is not entirely new in writing with technology: studies show that undergraduate writers planned less before writing when they used a word processor than with pen and paper (Hayes, 1996; Haas, 2013).
-If the shift from pen and paper to word processors reduced the (presumed) need for planning, then it is natural to assume that there will be bigger impacts with the shift from manual writing to LLMs.
-Now that the generation of text itself can be done effortlessly within seconds, we see planning to primarily emerge as a result of the human writer reflecting on the AI-generated drafts.
-It is not that the value of planning is diminishing, but that AI co-writers may encourage users to get something down on paper first, then revise and plan accordingly, nudging writing more towards an empirical “trial and error” process than a fully planned out one, similar to prototyping.
-Buschek’s work on comparing the current human-AI collaborative writing scene with literary collage from avant-garde literature (Buschek, 2024) frames this shift elegantly: writing with LLMs is closer to the assembling of “prefabricated text”, consisting of multiple voices and sources.
-Design implication #2: Equip users with granular revision control & Design novel methods to effectively prototype complex documents
+Interestingly, we are not entirely new to this shift: studies show that undergraduate writers planned less before writing when they used a word processor than with pen and paper (Hayes, 1996; Haas, 2013).
+If the shift from pen and paper to word processors reduced the (presumed) need for planning, then it is natural to assume that there will be an even larger reduction when shifting from manual writing to LLMs.
+That is not to say, however, that the value of planning is diminishing.
+Planning stays just as important, but from the standpoint of appropriate allocation of tasks (Vaccaro et al., 2024), it makes more sense for the AI to prioritize translation and reviews.
+Design implication #2: More rigor is required for helping human writers review&revise AI generated text
 Shifting the focus from planning to revision, however, also comes with a cost when it comes to LLMs.
 LLMs are prone to sycophancy bias (Sharma et al., 2023), essentially making it a “yes man” regardless of factual truth.
-If so, is it also unrealistic to expect genuine critical feedback from LLMs?
-Hints can be drawn from the Re3 framework (Ruan et al., 2024), which proposes the taxonomy of revision in Figure 3.
-By dividing revision by its qualitative dimensions such as action, intent and granularity, and further subdividing each by their characteristics, we could enable users to provide more granular instructions to LLMs on the level of desired revision, such as requesting for grammar or clarity related edits for intent on sentence or subsentence granularity, but forbidding factual content updates which LLMs are notoriously known to be unreliable with.
-Given that the revision has long been mistakenly seen as just the “tidying-up activity” after completing a first draft (Faigley and Witte, 1981), helping the human writers to get accustomed to these more granular terminologies around revision is then perhaps the key agenda here.
-We also acknowledge that the instantaneous and effortless text-generation can potentially lead to the loss of critical thinking for the human writers, as already seen with image generation tools (Wadinambiarachchi et al., 2024).
-If we are to blindly generate documents or emails to send to our peers without reviewing their output, then the results would certainly be catastrophic, both for our cognitive process, and for workplace collegiality and responsibility.
-The problem here is not the ease of text generation, but the difficulty in reviewing and iterating fully developed text.
-Perhaps the question we should be asking ourselves here is: how can we allow for effective prototyping of complex documents?
-For example, by leveraging the LLM’s semantic capabilities it is now possible for us to draft with smaller or summarized chunks of text, then move on to a fully fleshed document, similar to evolving prototypes from low-fidelity to high-fidelity.
-We see this to be a highly interesting and novel challenge that we welcome researchers to build further.
+If so, is it unrealistic to expect genuine critical feedback from LLMs?
+Hints can be drawn from the Re3 framework (Ruan et al., 2024), which proposes the taxonomy of revision as shown in Figure 4.
+By dividing revision by its qualitative dimensions such as action, intent and granularity, and further subdividing each by their characteristics, we could enable users to provide more granular control over revision instructions, such as requesting for grammar or clarity related edits for intent on sentence or subsentence granularity, but forbidding factual content updates which LLMs are notoriously known to be unreliable with.
+Given that the revision has long been mistakenly seen as just the “tidying-up activity” after completing a first draft (Faigley and Witte, 1981), helping the human writers to get accustomed to these more granular terminologies around revision is perhaps the key agenda here.
+In a similar vein, Laban et al.
+(Laban et al., 2024) proposed the Warn-Verify-Audit framework, where new information introduced to the document by LLMs are constantly flagged, and are automatically searched online or sent to an external auditor for manual fact checking.
+We see similar interesting fact-checking approaches emerging (Liu et al., 2024; Amirizaniani et al., 2024; Dietz, 2024), all of which are human-in-the-loop approaches to maximize the strengths of manual oversight and LLM automation.
+Such novel approaches to support human writes review and revise will eventually need to be built into human-AI collaborative writing tools, to draw out the best of human-AI performance.
 4.2.3. Group awareness is a key to success
@@
 Design implication #3: Provide a moment for writers to reflect on their expectations and needs from the AI co-writer
-While a direct application of human-human group awareness to human-AI collaboration could be considered excessive anthropomorphism, we believe that defining the human’s expectation for AI tool’s roles and boundaries prior to the writing could help mitigate under- and over-reliance, as seen in the field of Explainable AI.
+As stated in Section 1.1, we prefer to see AI simply as a tool and not a team mate.
+While the use of human-like metaphors are certainly of benefit in human-AI collaboration, anthropomorphizing AI as a team mate often misleads users to false expectations, which is a problem when AI simply cannot be accountable for written content (Stokel-Walker, 2023).
+Therefore we consider straightforward translations of human-human group awareness to human-AI collaboration (such as considering the AI’s motivations, emotional state or preferred writing styles) to be excessive anthropomorphism.
+However, the field of Explainable AI has also shown that uncovering the human’s expectation for AI tool’s roles and boundaries prior to its use could help mitigate under- and over-reliance.
 For instance, Cai et al (Cai et al., 2019b) found that, in clinical contexts, clinicians wished to know the training data, the medical views and the design objectives of the AI, far before engaging in the explanations of their diagnosis recommendations.
-As long as it is in service of improving the user’s motivation towards the task at hand, it may be worthwhile to provide a quick onboarding to align on the user’s expectations on the AI tools’s roles and capabilities prior to commencing the writing.
+As long as it is in service of improving the user’s motivation towards the task at hand, it may be worthwhile to provide a quick onboarding to align on the user’s expectations on the AI tool’s roles and capabilities, prior to commencing the writing.
 We see Wiethof et al.
 (Wiethof et al., 2021)’s work to be in line with this principle: they provided users with a written self-introduction by the AI agent prior to the writing, which led to “realistic expectations”.
-However, their implementation also came with the literal adoption of Nass and Moon’s Social Response Theory (Nass and Moon, 2000) such as providing social identities to the agent, which we consider unnecessary anthropomorphism.
-Similarly, Cress and Kimmerle (Cress and Kimmerle, 2023) suggest refining the human-human group awareness to suite the more asymmetrical human-AI relationship, and suggests the need for AIs to develop theory of mind, the cognitive ability to be aware of others’ mental states.
-A good balance here may be to simply provide a moment for the writer to individually reflect on what support they are seeking from the AI co-writer, both to nurture the self-awareness as well as to inform the AI tool of the user’s expectations and goals, but not introduce the AI co-writer as another team mate with its own values and writing styles.
+However, their implementation also came with a rather literal adoption of Nass and Moon’s Social Response Theory (Nass and Moon, 2000) by providing social identities to the agent, which we see to be unnecessary anthropomorphism.
+Similarly, Cress and Kimmerle (Cress and Kimmerle, 2023) suggest refining human-human group awareness to suite the more asymmetrical human-AI relationship, and propose AIs to develop theory of mind, the cognitive ability to be aware of others’ mental states.
+A good balance here may be to simply provide a moment for the writer to individually reflect on what support they are seeking from the AI co-writer, both to nurture their own self-awareness as well as to inform the AI tool of the user’s expectations and goals, but stay away from introducing the AI co-writer as another team mate with its own values and writing styles.
 When it comes to multiple human authors writing collaboratively with an AI co-writer, then it also becomes the AI co-writer’s role to facilitate team communication and at times resolve conflicts or consensus between the humans, which is a critical factor in the offloading of cognitive load in teamwork as we explain in 4.2.7.
@@
 Visual representation of the semantic document structure as seen in VISAR (Zhang et al., 2023) serves as a great inspiration here.
-Providing a visual cue of the current document structure would not just allow the user to be in constant alignment with the AI, but would essentially force the LLMs to engage in some form of “planning”, which would not otherwise be technically needed (Elkins, 2024) (unless specified to do so, such as with Chain-of-Thought prompting).
+Providing a visual cue of the current document structure would not just allow the user to be in constant alignment with the AI, but would also involve the LLMs to engage in a form of “planning”, which they do not require otherwise (Elkins, 2024) (unless specified to do so, such as with Chain-of-Thought prompting).
 Ultimately, it would enable the user to have an understanding of “how their text works” (Rahhal et al., 2007), which will be valuable regardless of the mode of collaboration.
@@
 If readers are a part of the writing, then it would mean that readers are inseparable from the LLM’s text generation process.
-A straightforward implementation would be to simply include the intended audience profile to the prompt, but we do not see this to be the solution.
+A straightforward implementation would be to simply include the intended audience profile to the prompt, but we do not see this to be a solution.
 The fundamental implication here is that writers should always be conscious of the reader’s grammar, language and existing knowledge to maximize the ease of comprehension (Rahhal et al., 2007).
@@
 (Dhillon et al., 2024) explored the effects of scaffolding during human-AI collaborative writing, and found that high scaffolding (i.e.
-next paragraph suggestion) resulted in higher writing quality and productivity, but also resulted in the decrease of text ownership and overall satisfaction.
-As more and more writing is conducted by AI, it is inevitable that our sense of ownership towards the text will decrease.
+suggestions for entire paragraphs) resulted in higher writing quality and productivity, but also resulted in the decrease of text ownership.
+As more and more writing is conducted by AI, it seems inevitable that our sense of ownership towards the text will decrease(Kosmyna et al., 2025).
 Since our scope for this paper is the pragmatic support of task completion, we do not necessarily see this as a problem.
-Having said that, the concept of authorship is directly tied with the author’s accountability (Stokel-Walker, 2023), and given that only the human writers can be kept accountable for the consequences of any human-AI collaborative writing (Stokel-Walker, 2023), design considerations are necessary to clarify where the accountability lies.
-Design implication #6: Design data structures and visual cues to distinguish human-written text from AI-written ones & Provide built-in verification methods for AI-generated text
+Having said that, the concept of authorship is directly tied with the author’s accountability (Stokel-Walker, 2023), and given that only the human writers can be kept accountable for the consequences of any human-AI collaborative writing (Stokel-Walker, 2023), design considerations to clarify where the accountability lies seems necessary.
+Design implication #6: Design data structures and visual cues to distinguish human-written text from AI-written ones
 We believe that keeping track of human contribution vs LLM contribution is a crucial aspect in future collaborative writing tools.
@@
 Such data structures would allow for collaborative writing tools to visually distinguish human written text from AI written ones, enhancing the notion of accountability for the user.
-Their work also proposes the Warn-Verify-Audit framework, where new information introduced to the document by LLMs are flagged, and can be automatically searched online or sent to an external auditor for manual fact checking.
-We see similar interesting fact-checking approaches emerging (Liu et al., 2024; Amirizaniani et al., 2024; Dietz, 2024), all of which are human-in-the-loop approaches to maximize the strengths of manual oversight and LLM automation.
-While these studies are beyond the scope of this paper, they would inevitably need to be eventually built into human-AI collaborative writing tools.
+In Kosmyna et al.
+(Kosmyna et al., 2025)’s study on essay writing with ChatGPT, they also touch on how lower ownership of leads to the “diminished sense of cognitive agency”.
+That is, when we feel that a written text is AI’s and not ours, we will feel less control over the product, leading to decrease in error monitoring.
+Perhaps then it is important to go beyond visualization of writers, and to provide interactions that will encourage human writers to increase their sense of ownership over the text.
 4.2.7. Collaborative writing tools serve as external cognition
@@
 This allows us to circle back to the very reason why human-AI collaboration is beneficial in work-context writing tasks in the first place: the primary purpose of involving LLMs in collaborative writing should be to get the job done in the most effective manner, and the offloading of cognitive resources is a key criteria in achieving this goal.
-As we design AI co-writer tools, we must take great care in making sure that the tool is not resulting in unwanted cognitive load, bringing more disruption than improvement.
-Unnecessary cognitive load can be witnessed at various moments throughout the writing process.
+As we design AI co-writer tools, we must take great care in making sure that the tool is not resulting in additional cognitive load, bringing more disruption than improvement.
+Undesired cognitive load can be witnessed at various moments throughout the writing process, not just in the writing.
 For example, Dhillon et al.
@@
 (Larsen-Ledet et al., 2020) shed light from a socio-technical perspective, finding that writers often switch between multiple writing tools, not just for functional reasons, but to get away from the collaborative mode to “sit for myself.” That is, writers sometimes prefer to stay away from collaboration to stay focused.
-Traditionally, cognitive load has been measured in HCI through post-task questionnaires, such as with the NASA task load index (Hart and Staveland, 1988), although recently the validated of this for HCI has been questioned (Babaei et al., 2025).
-Emerging frameworks such as the Overall Writing Effectiveness framework (Wilbers et al., 2024) could also serve as an alternative metric in surfacing the impacts that LLMs bring to collaborative writing.
 Design implication #7: Make sure that the AI co-writer is contributing to both the production of text as well as the facilitation of the process itself
@@
 This is inspired greatly by the way Fiore and Wiltshire (Fiore and Wiltshire, 2016) emphasize the importance of distinguishing teamwork (behaviors required for working together with others) and taskwork (functions required for getting the task done).
-Tools support teamwork through the visualization of group awareness, team progress and facilitating communication, while tools also support taskwork through the direct assistance of the task (e.g.
+Tools can support teamwork through the visualization of group awareness, team progress and facilitating communication, while tools can also support taskwork through the direct assistance of the task (e.g.
 writing or decision making).
@@
 5. Discussion
+5.1. Applying existing human-human collaborative writing theories/frameworks to the human-AI paradigm
 We now come back to our initial research question: what insights can we draw from theories/frameworks on human-human collaborative writing to apply to the design of human-AI writing tools?
@@
 We outline our views in Figure 11.
-While the role of humans undeniably changes from a hands-on writer to more of an editorial decision maker (Buschek, 2024), we believe that the cognitive process of writing will not and should not change — at least for good writing.
-This is to make sure that the writer continues to stay accountable for the writing they produce (which the AI cannot).
-Because of this, the non-linear, interchangeable nature of writing that studies based on the Cognitive Process Theory of Writing (Flower and Hayes, 1981) will stay as a core foundation of future collaborative writing tool designs.
-However, because the speed and ease of translating ideas to words will increase dramatically, we believe the writing process to become closer to an empirical “trial and error” process analogous to prototyping, resulting in less need for fully fledged planning, and more support needed for rigorous revision methodologies.
-In particular, tools will need to equip writers with enhanced understanding and terminologies for different revision types, to draw more granular boundaries for the AI support.
-This shift in turn would further elevate the importance of coherence support in human-AI collaborative writing, which has been an ongoing challenge in this field to this day.
-Given the unprecedented semantic classification capabilities of LLMs, we are optimistic about the impact that LLMs can bring to the alignment of document structures through real-time visualization, which can positively impact not just the human-AI collaborative writing but human-human writing as well.
-Since we take the stance that AIs are no more than a tool and humans should always remain in control, teamwork-related factors such as group awareness, consensus building and authorship, which have been much discussed in collaborative writing thus far, will not apply in the human-AI paradigm, except between multiple human co-authors.
+While the role of humans will change from a hands-on writer to more of an editorial decision maker (Buschek, 2024), we believe that the cognitive process of writing will not and should not change — at least for good writing.
+This is to make sure that the writer continues to stay accountable for the writing they produce (which the AI cannot) (Stokel-Walker, 2023).
+Because of this, the non-linear, interchangeable nature of writing based on the Cognitive Process Theory of Writing (Flower and Hayes, 1981) should stay as a core foundation of future collaborative writing tool designs.
+However, as more humans choose to delegate the actual writing to the essentially effortless AI generation, the legacy planning/translating/reviewing process shifts to a planning/waiting/reviewing process, thereby making the human writer’s cognitive process discontinuous due to the waiting that occurs in between.
+It becomes crucial then for the collaborative writing tool to provide support in streamlining the otherwise broken cognitive process of the human writer.
+For this we suggest a “prototyping” approach, where the tool works to shorten the waiting stage as well as accelerate the overall cognitive process, to allow human writers to test new ideas at a similar cognitive speed as when they are writing themselves.
+For instance, tools could enable quick iterations of the cognitive process by starting with digestible, smaller chunks of text, and gradually moving on to a fully fleshed-out document, much like evolving prototypes from low-fidelity to high-fidelity.
+Consequently, more rigor will be required for novel revision methodologies and interactions to allow for human writers to quickly read, review and edit AI generated text.
+Tools will need to equip writers with enhanced understanding and terminologies for various revision types, to set more granular boundaries for the desired AI support and to draw out the best of human and AI performance.
+As AI takes on more of the content writing, the ongoing challenge of coherence support will be of further importance.
+Given that LLMs provide a level of semantic generalization and abstraction not achievable with prior AI models, we are optimistic about the impact that LLMs can bring to the alignment of document structures, such as through real-time visualization.
+These interventions will positively impact not just the human-AI collaborative writing but human-human writing as well.
+Since we take the stance that AIs are no more than a tool and humans should always remain in control, teamwork-related factors such as group awareness, consensus building and authorship, which have been much discussed in collaborative writing studies thus far, will not apply in the human-AI paradigm, except between multiple human co-authors.
 Therefore, when it comes to the writing strategies, it seems inevitable that most human-AI collaborative writing will end up as single-author writing, or will be conducted in the hierarchical mode where the AI co-writer will primarily translate ideas to words instantaneously under human supervision, and the humans will primarily take an editorial role.
 The more equally distributed forms of collaborative writing, known as reactive writing, parallel writing or the dialogic mode, will remain a strategy unique to human collaboration.
-However, given that the reduction of cognitive load results not just from the direct support of text production but in streamlining the collaboration process as well, it is an equally important role of the AI co-writer to remove any unnecessary frictions in the collaboration process itself, particularly when there are more than one human writers involved.
+However, given that the reduction of cognitive load derives not just from the direct support of text production but from the streamlining of the overall collaboration process, it is equally important that the AI co-writer works to remove any unnecessary frictions in the collaboration process itself, particularly when there are more than one human writers involved.
 Finally, the notion that writing is socially constructed through readers remains to be just as valid, and we urge researchers to explore the implications that LLM’s diverse text generation capabilities will have to this space.
 At the same time, we acknowledge that this will only apply while humans are the readers: if human and AIs are co-writing, there is no reason to doubt that humans and AIs will co-read, and if (willingly or unwillingly) documents are to be agentically processed in the future, this will signal another significant turning point for writing in general.
+Much like how word-processors had a profound impact on the way that we wrote with pen and paper, LLMs seem to be fundamentally shifting the way that we approach writing.
+While there are certainly signs of de-skilling, as HCI researchers we are also curious to see what synergistic effects arise from this shift, resulting in novel mindsets and methodologies of writing.
+As Brinkmann et al.
+(Brinkmann et al., 2023) argue through the lens of “machine cultures”, our culture of writing is already being reshaped by our interactions with LLMs.
+5.2. Prompting as a form of writing
+In Figure 1, we assumed that, when writing with AI, prompting becomes the key component of the planning process.
+When prompting LLMs, we consider not just the content of the writing but the audience, length, structure, tone, and internal knowledge that we posses (often referred to as “context” in prompt engineering) that is required for the writing.
+This is much in line with how Flower and Hayes describe their model: they use the term planning in a “much broader sense”, where writers form abstract “representations” of ideas and goals, often involving retrieval from long-term memory (Flower and Hayes, 1981).
+The way they explain the iterative nature of goal-setting is also highly analogous to the way LLMs are used today, refining instructions through multiple trial and errors (Arawjo et al., 2023; Liu et al., 2022).
+However, it is important to note that prompting itself is also a form of writing.
+That is, the act of writing a prompt is also the very result of the writer’s cognitive writing process.
+In this scenario, the user would first engage in the planning/translating/reviewing to write a prompt, which would then be fed to an LLM for the user to engage this time in a planning/waiting/reviewing process.
+Further, if the writer chooses to use meta-prompting (Reynolds and McDonell, 2021) (using LLMs to create and refine prompts), then they go through the planning/translating/reviewing process to write a prompt that retrieves a prompt which they would revise and iterate through the planning/waiting/reviewing process, then apply the generated prompt to the LLM to generate the final outcome which again would be revised through the planning/waiting/reviewing process, and so forth.
+In either case, when viewed from the lens of this study, prompt writing is distinctly different from most forms of writing in the sense that it is a uniquely reader specific form of writing –in this case the reader of course being the LLM.
+Grammar or spelling errors are not rigorously checked, since they do not matter as long as the instruction gets through.
+Additional contexts are also often added without much thought, regardless of knowing whether the outcome would be improved from doing so(Liu et al., 2022).
+Prompt engineering or defense techniques (Liu et al., 2023) (such as the ”sandwich prevention”, which refers to repeating the certain instructions at the beginning and the end of a prompt) are often highly unnatural forms of writing.
+Despite these peculiarities as a form of writing, it is also interesting that the concept of authorship seems to apply to prompts: platforms such as PromptBase444See https://promptbase.com/ allow users to sell and buy effective prompts.
+As such, prompting itself can be seen as the act of writing, with distinct characteristics.
+In this sense, when humans co-write with AI, the process begins with a legacy human mode of writing, resulting in a mix of both approaches that we outlined in Figure 11.
+How does this affect the overall process of writing with AI?
+Is it necessary to intentionally separate the initial prompting (i.e.
+the legacy planning/translating/reviewing process) from the human-AI mode of writing (i.e.
+the planning/waiting/reviewing process) that follows?
+These are still early questions in trying to understand how LLMs are affecting the way that we write, and we welcome more researchers to explore this space.
 6. Limitations
-Finally, we would like to acknowledge some general limitations of our literature analysis approach.
+Finally, we acknowledge some general limitations of our literature analysis approach.
 First, we chose to look at Web of Science and the ACM Digital Library, but it is possible that we are missing collaborative writing theories or frameworks that were not covered by these two databases.
+For example, among the 6 articles that we added through backward referencing, 4 were found to be in the ProQuest database (the remaining 2 were the cognitive process theories by Flower and Hayes (Flower and Hayes, 1981; Hayes, 1996), which are fundamental to this study but are technically not collaborative writing).
 We are, however, confident that we have sufficiently covered at least the most influential ones commonly found in past literature.
@@
 7. Conclusion
-In this paper, we have argued for the need to reassess foundational theories and frameworks of collaborative writing to adapt to the human-AI paradigm, as LLMs increasingly become our daily co-writers in a uniquely asymmetrical relationship unlike any we have previously encountered.
+In this paper, we have argued for the need to reassess foundational theories and frameworks of collaborative writing to adapt to the human-AI paradigm, as LLMs increasingly become our daily co-writers.
 We conducted a critical review of existing theories and frameworks on human-human collaborative writing to assess their relevance to the human-AI paradigm, and drew seven insights and implications that can be applied to the design of human-AI writing tools.
-We found that, as LLMs nudge the writing process more towards an empirical “trial and error” process analogous to prototyping, the non-linear cognitive process of writing will stay the same, but more attention will be required for rigorous revision methodologies.
+We found that, as we delegate more writing to AI, our cognitive process shifts from the traditional planning/translating/reviewing process to a planning/waiting/reviewing process, breaking the process due to the waiting that occurs in between.
+To ensure that our cognitive process remains intact, we suggest a “prototyping” approach, in which the tool allows for quick iterations of the cognitive process by starting with smaller or summarized chunks of text, and gradually moving on to a fully fleshed document.
+For this we would need to come up with novel revision methodologies and interactions.
 In turn, this may reduce the focus on the planning aspects of collaborative writing, further increasing the importance of coherence support between human and AI writing.
-We believe that the LLM’s unprecedented semantic capabilities can bring a break through to this ongoing challenge, through real time visualizations of the document’s semantic structure.
+We believe that the LLM’s strong semantic capabilities can provide analyses that were previously difficult with traditional methods, such as real time visualizations of a document’s semantic structure.
 Teamwork-related factors such as group awareness, consensus building and authorship, which have been central discussions in human-human collaborative writing, will not apply to the human-AI paradigm, resulting in the majority of human-AI collaborative writing being categorized in the single-author writing strategy.
@@
 Acknowledgements.
-The interface designs in this paper used Google’s Material Icons444See https://fonts.google.com/icons provided under the Apache License, Version 2.0.References
+The interface designs in this paper used Google’s Material Icons555See https://fonts.google.com/icons provided under the Apache License, Version 2.0.References
 - (1)
@@
 2024).
-- Babaei et al.
-(2025) Ebrahim Babaei, Tilman Dingler, Benjamin Tag, and Eduardo Velloso.
-2025.
-Should we use the NASA-TLX in HCI?
-A review of theoretical and methodological issues around Mental Workload Measurement.
-International journal of human-computer studies 201, 103515 (June 2025), 103515.
-doi:10.1016/j.ijhcs.2025.103515
+- Arawjo et al.
+(2023) Ian Arawjo, Priyan Vaithilingam, Martin Wattenberg, and Elena Glassman.
+2023.
+ChainForge: An open-source visual programming environment for prompt engineering.
+In Adjunct Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology.
+ACM, New York, NY, USA.
+- Brinkmann et al.
+(2023) Levin Brinkmann, Fabian Baumann, Jean-François Bonnefon, Maxime Derex, Thomas F Müller, Anne-Marie Nussberger, Agnieszka Czaplicka, Alberto Acerbi, Thomas L Griffiths, Joseph Henrich, Joel Z Leibo, Richard McElreath, Pierre-Yves Oudeyer, Jonathan Stray, and Iyad Rahwan.
+2023.
+Machine culture.
+Nat.
+Hum.
+Behav.
+7, 11 (Nov.
+2023), 1855–1868.
 - Buschek (2024) Daniel Buschek. 2024. Collage is the new writing: Exploring the fragmentation of text and user interfaces in AI tools. arXiv [cs.HC] (May 2024).
@@
 2025).
-- Hart and Staveland (1988) Sandra G Hart and Lowell E Staveland.
-1988.
-Development of NASA-TLX (task load index): Results of empirical and theoretical research.
-In Advances in Psychology.
-Advances in psychology, Vol.
-52.
-Elsevier, 139–183.
 - Hayes (1996) J Hayes. 1996. A new framework for understanding cognition and affect in writing. The science of writing: Theories, methods, individual differences, and applications. 432 (1996), 1–27.
@@
 2025).
+- Kosmyna et al.
+(2025) Nataliya Kosmyna, Eugene Hauptmann, Ye Tong Yuan, Jessica Situ, Xian-Hao Liao, Ashly Vivian Beresnitzky, Iris Braunstein, and Pattie Maes.
+2025.
+Your brain on ChatGPT: Accumulation of cognitive debt when using an AI assistant for essay writing task.
+arXiv [cs.AI] (June 2025).
 - Laban et al.
@@
 - Lay and Karis (1991) Mary M Lay and William M Karis. 1991. Collaborative Writing in Industry: Investigations in Theory and Practice. Baywood Press.
+- Lee et al.
+(2025) Hao-Ping (hank) Lee, Advait Sarkar, Lev Tankelevitch, Ian Drosos, Sean Rintel, Richard Banks, and Nicholas Wilson.
+2025.
+The impact of generative AI on critical thinking: Self-reported reductions in cognitive effort and confidence effects from a survey of knowledge workers.
+In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems.
+ACM, New York, NY, USA, 1–22.
 - Lee et al.
@@
 2024), 1–44.
+- Liu et al.
+(2022) Jiachang Liu, Dinghan Shen, Yizhe Zhang, Bill Dolan, Lawrence Carin, and Weizhu Chen.
+2022.
+What makes good in-context examples for GPT-3?.
+In Proceedings of Deep Learning Inside Out (DeeLIO 2022): The 3rd Workshop on Knowledge Extraction and Integration for Deep Learning Architectures.
+Association for Computational Linguistics, Stroudsburg, PA, USA, 100–114.
+- Liu et al. (2023) Yupei Liu, Yuqi Jia, Runpeng Geng, Jinyuan Jia, and Neil Zhenqiang Gong. 2023. Formalizing and benchmarking prompt injection attacks and defenses. arXiv [cs.CR] (Oct. 2023).
 - Lowry et al.
@@
 - Lunsford and Ede (1986) A Lunsford and Lisa S Ede. 1986. Why write…Together: A research update. Rhetoric Review 5 (Sept. 1986), 71–77.
+- Ma et al.
+(2024) Shuai Ma, Qiaoyi Chen, Xinru Wang, Chengbo Zheng, Zhenhui Peng, Ming Yin, and Xiaojuan Ma.
+2024.
+Towards human-AI deliberation: Design and evaluation of LLM-empowered deliberative AI for AI-assisted decision-making.
+arXiv [cs.HC] (March 2024).
 - Mendoza-Chapa et al.
@@
 BMJ 339, jul21 1 (July 2009), b2535.
+- Naikar et al.
+(2025) Neelam Naikar, Robert Hoffman, Emilie M Roth, Gary Klein, Laura G Militello, and Cindy Dominguez.
+2025.
+Should we Make AI More Tool-like or Teammate-Like?
+J.
+Cogn.
+Eng.
+Decis.
+Mak.
+15553434251346904 (June 2025).
 - Nass and Moon (2000) Clifford Nass and Youngme Moon. 2000. Machines and mindlessness: Social responses to computers. J. Soc. Issues 56, 1 (Jan. 2000), 81–103.
@@
 IEEE, 484–493.
+- Reynolds and McDonell (2021) Laria Reynolds and Kyle McDonell.
+2021.
+Prompt programming for large language models: Beyond the few-shot paradigm.
+In Extended Abstracts of the 2021 CHI Conference on Human Factors in Computing Systems.
+ACM, New York, NY, USA.
 - Ruan et al.
@@
 arXiv [cs.HC] (May 2024).
-- Wang et al.
-(2020) Dakuo Wang, Elizabeth Churchill, Pattie Maes, Xiangmin Fan, Ben Shneiderman, Yuanchun Shi, and Qianying Wang.
-2020.
-From human-human collaboration to human-AI collaboration.
-In Extended Abstracts of the 2020 CHI Conference on Human Factors in Computing Systems.
-ACM, New York, NY, USA.
 - Weld and Bansal (2019) Daniel S Weld and Gagan Bansal. 2019. The challenge of crafting intelligible intelligence. Commun. ACM 62, 6 (May 2019), 70–79.
@@
 ACM, New York, NY, USA, 1–30.
+- Zhu et al.
+(2024) Tiffany Zhu, Iain Weissburg, Kexun Zhang, and William Yang Wang.
+2024.
+Human bias in the face of AI: The role of human judgement in AI generated text evaluation.
+arXiv [cs.CL] (Sept.
+2024).
 Appendix A Appendix: Systematic Review Screening Questions